### v2

# Testing Monotonicity of Real-Valued Functions on DAGs

Abstract
We study monotonicity testing of real-valued functions on directed acyclic graphs (DAGs) with $n$ vertices. For every constant $\delta>0$, we prove a $\Omega(n^{1/2-\delta}/\sqrt{\varepsilon})$ lower bound against non-adaptive two-sided testers, nearly matching the classical $O(\sqrt{n/\varepsilon})$-query upper bound. For constant $\varepsilon$, we also prove an $\Omega(\sqrt{n})$ lower bound for adaptive one-sided testers, whereas previously only an $\Omega(\log n)$ lower bound was known. Both lower bounds rely on positive-matching Ruzsa–Szemerédi families. On the algorithmic side, we give simple non-adaptive one-sided testers with query complexity $O(\sqrt{m\,\ell}/(\varepsilon n))$ and $O(m^{1/3}/\varepsilon^{2/3})$, where $m$ is the number of edges in the transitive reduction and $\ell$ is the number of edges in the transitive closure. For constant $\varepsilon>0$, these improve over the previous $O(\sqrt{n/\varepsilon})$ bound when $m\ell=o(n^{3})$ and $m=o(n^{3/2})$, respectively.
1 Introduction
Testing monotonicity on structured domains is a central problem in property testing and a standard example in the sublinear-time model; see, e.g., [17, 23, 8, 22, 29]. It was first studied on the Boolean hypercube and product domains and has since been extended to richer product structures and general posets/graph families [2, 7, 11, 16, 19, 21, 25, 28]. One way to model posets is via reachability in directed acyclic graphs (DAGs). In this work we study the massively parameterized model [29, 23], where the entire DAG is given explicitly.
1.1 Problem and background
Let $G=(V,E)$ be a DAG with $n=|V|$ vertices. A function $f:V\to\mathbb{R}$ is monotone if $f(u)\leq f(v)$ whenever $(u,v)\in E$. Given full access to $G$, a proximity parameter $\varepsilon>0$, and oracle access to $f$, the goal is to distinguish monotone functions from those that are $\varepsilon$-far from monotonicity with constant success probability. We do not charge for preprocessing or computations on $G$; only queries to $f$ are counted. We consider both non-adaptive and adaptive testers, with one-sided or two-sided error. A tester is non-adaptive if all queries are fixed in advance (adaptive otherwise), and one-sided if it always accepts monotone functions (two-sided otherwise). Much of the monotonicity-testing literature on both upper and lower bounds focuses on product domains such as the hypercube and hypergrid, where in some settings tight bounds are known [5, 10, 11, 12, 13, 14, 15, 16].
For general posets, the classical non-adaptive one-sided tester of Fischer et al. [19] achieves
$O\!\left(\min\left\{\sqrt{\frac{n}{\varepsilon}},\ \frac{\ell}{\varepsilon n}\right\}\right)$
queries, where $\ell$ is the number of comparable pairs, i.e., the number of edges in the transitive closure of $G$. They also show non-adaptive two-sided lower bounds, for Boolean ranges, based on Ruzsa–Szemerédi-type constructions [19], yielding a bound of $n^{\Omega(1/\log\log n)}$. This is still far from the abovementioned upper bound, leaving a large gap.
1.2 Our results
Our main contribution is improved lower and upper bounds for the query complexity of monotonicity testing on DAGs: we strengthen lower bounds in both non-adaptive and adaptive models, and complement them with improved non-adaptive one-sided upper bounds in sparse regimes. The quantitative statements and their comparison with prior work are summarized in Table 1.
Table 1: Comparison of the main bounds with prior work.
Model Previous best This work
Lower bounds
Non-adaptive two-sided $n^{\Omega(1/\log\log n)}$ [19] $\Omega(n^{1/2-\delta}/\sqrt{\varepsilon})$ (Theorem 3.12)
Adaptive one-sided $\Omega(\log n)$ [17] $\Omega(\sqrt{n})$ (Theorem 4.28)
Upper bounds
Non-adaptive one-sided $O\!\left(\min\{\sqrt{n/\varepsilon},\,\ell/(\varepsilon n)\}\right)$ [19] $O(\sqrt{m\,\ell}/(\varepsilon n))$ (Theorem 5.1), $O(m^{1/3}/\varepsilon^{2/3})$ (Theorem 5.2)
Lower bounds and interpretation.
In the non-adaptive regime, the classical lower bound for general posets is $n^{\Omega(1/\log\log n)}$ [19], far from the $\sqrt{n/\varepsilon}$-type upper bound. For every constant $\delta>0$ and infinitely many $n$, we construct $n$-vertex DAGs for which every non-adaptive two-sided $\varepsilon$-tester needs $\Omega(n^{1/2-\delta}/\sqrt{\varepsilon})$ queries (Theorem 3.12), near $\sqrt{n/\varepsilon}$ up to $n^{o(1)}$ factors. We also briefly discuss bounded ordered ranges: for finite ordered ranges of size $(\log n)^{C}$ with $C>2$, we prove a polynomial non-adaptive two-sided lower bound $\Omega(n^{1/2-1/C-o(1)}/\sqrt{\varepsilon})$ (Section 3.6).
In the adaptive regime, logarithmic lower bounds are known already on the total order (equivalently, path graph) [17]. For constant $\varepsilon$ and infinitely many $n$, we prove an $\Omega(\sqrt{n})$ lower bound on bipartite $n$-vertex DAGs for adaptive one-sided testers (Theorem 4.28).
Upper bounds.
We next outline upper bounds that leverage sparsity of the reachability structure. We write $m$ for the number of edges in the input DAG and $\ell$ for the number of comparable pairs. See Figure 1 for a comparison of the resulting regimes.
  • A bound obtained by combining sampling from the transitive reduction and transitive closure: there is a tester with $q=O\!\left(\sqrt{m\,\ell}/(\varepsilon n)\right)$ queries (Theorem 5.1).
  • An $\ell$-independent bound: there is a tester with $q=O\!\left(m^{1/3}/\varepsilon^{2/3}\right)$ queries (Theorem 5.2).
These upper bounds can improve the $O(\min\{\sqrt{n/\varepsilon},\ell/(\varepsilon n)\})$ bound in sparse regimes, complementing the lower bounds above. They explicitly exploit both the transitive reduction and the transitive closure, as in work on reachability sparsification and TC-spanner-based testers and reconstruction on posets [6, 7, 28, 31].
Figure 1: Comparison of query complexities for monotonicity testing, where $m=\Theta(n^{c})$ and $\ell=\Theta(n^{d})$ and $\epsilon>0$ is a constant. If the underlying undirected graph is connected, then $n\leq m\leq n^{2}$ so $1\leq c\leq 2$, and since $m\leq\ell\leq n^{2}$ we have $c\leq d\leq 2$. Each colored region indicates which query complexity achieves the lowest value. Red: $O(\sqrt{m\ell}/(\varepsilon n))$ (Theorem 5.1), Blue: $O(m^{1/3}/\varepsilon^{2/3})$ (Theorem 5.2), Yellow: $O(\sqrt{n/\varepsilon})$ ([19]). The additional bound $O(\ell/(\varepsilon n))$ is omitted since it is never optimal in this range.
1.3 Technical overview
1.3.1 Non-adaptive two-sided lower bound
Lower bounds: from induced matchings to positive matchings.
The classical lower-bound approach for monotonicity testing on general posets, pioneered by Fischer et al. [18], is built around Ruzsa–Szemerédi (RS) graphs: one packs many large induced matchings and uses them to generate hard YES/NO instances. The relevance of matchings is that violated pairs in a matching are vertex-disjoint, so repairing all of them requires changing at least one endpoint per pair, which directly yields a distance-from-monotonicity lower bound. For query lower bounds, matchings also localize evidence: a tester typically needs to query both endpoints of one matched pair to observe informative violations. Inducedness then plays a dual role there. It is what makes the violation pattern “clean” for Boolean-valued functions (extra comparabilities among matching endpoints would immediately create additional violations), and at the same time it becomes the bottleneck: extremal graph theory severely limits how many linear-size induced matchings can coexist in a single graph, which is why RS-based arguments stall far below the $\tilde{\Theta}(\sqrt{n})$ regime.
Our first step is to identify the right analogue of induced matchings for real-valued functions. On bipartite DAGs (all edges oriented left-to-right), the question “can a matching be realized as exactly the set of violated edges?” has a precise answer: this holds if and only if the matching is positive (Section 3.1). Informally, “positive” means there is a vertex weight assignment $w$ such that $w(\ell)+w(r)>0$ exactly on matching edges and $w(\ell)+w(r)\leq 0$ on all other edges. Equivalently, positivity excludes alternating cycles/closed walks on the matched vertices (Theorem 2.3). This motivates the notion of positive-matching RS (PMRS) families, obtained by relaxing RS from induced matchings to positive matchings (Section 3.2). This relaxation is specific to the real-valued setting and bypasses the inducedness barrier in the Fischer et al. argument.
From PMRS to a near-tight non-adaptive lower bound.
Once a PMRS decomposition is available, the lower bound becomes essentially information-theoretic. Given $s$ edge-disjoint positive matchings $M_{1},\dots,M_{s}$ (each of linear size), we hide a random index $i\in[s]$ and construct two distributions over functions. The YES distribution is monotone, built from a base monotone function with a constant margin on non-matching edges (so small random noise does not create new violations), while the NO distribution is $\varepsilon$-far because it violates a constant fraction of edges in the hidden matching $M_{i}$ (Section 3.3). Crucially, the two distributions have identical single-vertex marginals; they differ only in a local correlation between the two endpoints of a matching edge. Therefore, a non-adaptive algorithm can obtain distinguishing information only if it queries both endpoints of some edge in $M_{i}$. Since the matchings are edge-disjoint, a $q$-query set can fully cover at most $O(q^{2})$ edges, and the probability of hitting the hidden matching index is at most $O(q^{2}/s)$, yielding the $\Omega(\sqrt{s})$ lower bound (Theorem 3.8). A useful technical fact (as with induced matchings) is that positivity is closed under taking submatchings (Section 3.1), which lets us refine each large matching into many smaller positive matchings and obtain the $1/\sqrt{\varepsilon}$ dependence (Sections 3.5 and 3.5).
Constructing PMRS: a polynomial-size packing beyond RS limitations.
The remaining challenge is to build explicit graphs with many large edge-disjoint positive matchings. This is precisely where the departure from induced matchings matters: unlike RS graphs, PMRS packings exhibit different extremal behavior and admit simple explicit constructions. Our construction (Section 3.4) uses vertex labels $(x,z)$ with $x\in[N]^{k}$ and an additional coordinate $z\in[N^{2}]$. Each matching is indexed by a shift vector $a$ and consists of translated edges $(x,z)\mapsto(x+a,\,z+\|a\|_{2}^{2})$. The nontrivial point is positivity: we design a weight function $w_{a}$ so that for an edge belonging to shift $b$, the positivity-witness sum $w_{a}(\ell)+w_{a}(r)$ (for edge endpoints $\ell,r$) collapses to a quantity of the form $1/2-\|b-a\|_{2}^{2}$, which is positive exactly when $b=a$ (Theorem 3.9). This quadratic construction separates the shift $a$ from every other shift through the sign of this sum and yields $s=\Omega(n^{k/(k+2)})$ edge-disjoint linear-size positive matchings. Combined with the PMRS $\Rightarrow$ lower-bound argument above, this yields our near-$\sqrt{n/\varepsilon}$ non-adaptive two-sided lower bound.
1.3.2 Adaptive one-sided lower bound
Violations on a hidden matching and the witness principle.
Moving from non-adaptive to adaptive testers, and from two-sided to one-sided error, requires a different lower-bound argument. In the bipartite-DAG setting (all edges oriented left-to-right), one-sidedness comes with a very strong “witness principle” (Section 4): if the tester has not queried a violated edge, then the partial transcript can always be extended to a globally monotone function. In other words, a one-sided tester is not “estimating distance” indirectly; it is literally searching for a witnessed violation. The challenge is that adaptivity seems to help precisely with such search problems: after observing a few values, the tester can steer future queries toward edges that are more likely to be violated.
Section 4 turns this search problem into an indexing problem. As in the PMRS-based non-adaptive lower bound, we start from a base graph containing many edge-disjoint positive matchings $M_{1},\dots,M_{s}$, and we hide a random index $I\in[s]$. The goal is to generate a distribution over functions such that (i) with high probability the sampled function is $\varepsilon$-far from monotone because it violates almost all edges of the hidden matching $M_{I}$, (ii) edges outside $M_{I}$ are almost never violated (so the tester cannot reject without learning something about $I$), and (iii) even an adaptive algorithm learns only limited information about $I$ per query. The last two requirements are exactly where naive constructions that simply prescribe a violating matching break down: once we condition on an adaptive prefix, the remaining unqueried variables are no longer independent, and a tester might try to create or expose violations on non-matching edges by strategically revealing endpoints.
Our lower-bound argument uses a truncated strongly log-concave prior, namely a Gibbs distribution $\{\mu_{i}\}_{i\in[s]}$ over real-valued labelings $f$ supported on a box. The Hamiltonian has three roles. A quadratic base term
$\frac{\alpha}{2}\sum_{v}f(v)^{2}+\lambda\sum_{(\ell,r)\in E}(f(\ell)-f(r))^{2}$
enforces global strong convexity, which later yields stability of conditional marginals under arbitrary transcript conditioning. On top of this, hinge-square potentials implement a monotonicity pattern with margin $\gamma$: for edges $e\notin M_{i}$ we add a penalty $\beta(f(\ell)-f(r)+\gamma)_{+}^{2}$, which favors $f(\ell)-f(r)\leq-\gamma$ (monotone with slack), while for edges $e\in M_{i}$ we add $\beta(\gamma-(f(\ell)-f(r)))_{+}^{2}$, which favors $f(\ell)-f(r)\geq\gamma$ (a violation with margin $\gamma$). Finally, we truncate to a box $B,B]^{V}$ so that the Lipschitz and concentration bounds remain uniform. Under this Gibbs distribution, edges of the hidden matching are violated with overwhelming probability, while edges outside the hidden matching are rarely violated, and this remains true under conditioning on adaptively revealed coordinates.
The rest of the proof has two technical parts: controlling spurious violations and bounding information leakage.
Spurious violations under adaptivity.
Even if a fresh non-matching edge is very unlikely to be violated unconditionally, an adaptive tester might try to bias it by first querying one endpoint and then deciding whether to query the other. The analysis therefore must be prefix-conditioned: after any transcript prefix, when the tester “closes” a new edge, the conditional probability of witnessing a violation on an edge outside $M_{I}$ must still be tiny. This is handled by a careful one-dimensional reduction on the difference coordinate $d=f(\ell)-f(r)$ for an edge $e=(\ell,r)$: after conditioning on everything except the two endpoints, the induced density of $d$ is log-concave with explicit strong curvature coming from the quadratic term. The hinge term sets a target point ($\gamma/2$ for matching edges and $\gamma/2$ for non-matching edges), and strong convexity translates “the derivative at the target has the right sign” into an exponentially small tail probability of landing on the wrong side.
The subtlety is the external drift contributed by all other incident constraints: under adaptive conditioning, this drift is random and depends on the revealed neighborhood. Section 4 controls this drift at the two critical points $d=\pm\gamma/2$ with a high-probability argument: for $C_{4}$-free base graphs, the relevant neighborhood variables become independent after conditioning on the endpoints, and strong log-concavity gives subgaussian tails for the drift; see Section 4.2. This is paired with a separate bound on the conditional mean of the drift; in the final instantiation on our explicit dense base graph (Section 4.7), this is the main analytic difficulty.
There is a second obstacle specific to the one-sided setting: the box truncation introduces boundary effects, and an adaptive tester might try to push queried values near $\pm B$, where the feasible interval of $d$ shrinks and the clean tail bound could fail. The fix is a deterministic inward bias (Sections 4.1 and 4.1): when $\beta>\lambda$, the hinge potentials create a deterministic force that pushes single-site conditionals away from the dangerous boundary strips. This yields a bound uniform over adaptive strategies on the probability that any query ever lands near $\pm B$ (Section 4.3), and hence bounds the probability that an adaptive strategy can exploit truncation to manufacture spurious witnesses. Putting these together yields a theorem controlling spurious queried violations (Section 4.3): with high probability, any $q$-query adaptive strategy witnesses no violated edge outside the hidden matching unless it hits an exponentially unlikely boundary event.
Information leakage from adaptive queries.
Even if spurious violations are suppressed, an adaptive tester could hope to learn the hidden index $I$ from the queried values, and then target an edge of $M_{I}$. Section 4 controls this via a mutual-information upper bound $I(I;T)$, where $T$ is the full adaptive transcript. The main analytic ingredient is a “Green-function comparison” inequality for uniformly convex Gibbs measures (Section 4.4): for any prefix conditioning $\tau$, the KL divergence between the one-site conditional distributions under $\mu_{i}(\cdot\mid\tau)$ and $\mu_{j}(\cdot\mid\tau)$ is bounded by a weighted sum of local discrepancies, with weights given by entries of $(\alpha I+2\lambda L_{\tau})^{-1}$, where $L_{\tau}$ is the Laplacian on the still-unqueried vertices after conditioning on the prefix $\tau$. This comparison needs no global weak-dependence or uniqueness assumption: it uses only the strong convexity provided by the quadratic term and survives arbitrary adaptive conditioning. A simple row-sum bound on this Green function then gives a per-query leakage bound $\kappa_{Q}$, leading to the budgeted bound $I(I;T)\leq q\kappa_{Q}$ uniformly over all adaptive strategies with $q\leq Q$ (Section 4.4).
With control of spurious queried violations and leakage in place, the lower bound becomes information-theoretic and well suited to one-sided testing. Because the matchings are edge-disjoint, any witnessed violated edge in $\bigcup_{i}M_{i}$ uniquely identifies the hidden index $I$. By the one-sided witness principle, rejection implies the transcript contains a witnessed violated edge; by Section 4.3, with high probability this violated edge lies in $M_{I}$. Thus any successful tester yields a decoder that recovers $I$ from $T$ with constant probability, and a standard Fano argument (Section 4.5) forces $I(I;T)\gtrsim\log s$. Comparing $\log s$ to the leakage upper bound $q\kappa_{Q}$ yields a lower bound for testers with query budget $q\leq Q$ (Theorem 4.22): below a certain budget threshold, no adaptive one-sided tester can simultaneously (a) keep spurious witnesses rare and (b) accumulate enough information to find a true matching-edge violation reliably.
Dense-graph instantiation and final bound.
Section 4.7 applies this argument to an explicit dense base graph: a $C_{4}$-free, almost-regular subgraph of the shift-matching construction (Section 4.7) with $s=\Theta(\sqrt{n})$ disjoint positive matchings. In this regime the auxiliary probabilities (matching-edge non-violation, spurious violations, boundary failures, and mean-drift failures) can all be driven superpolynomially small by setting the margin $\gamma$ to grow like $\sqrt{\log Q}$, while the per-query leakage remains on the order of $(\log Q)/\sqrt{n}$. Choosing $Q\asymp\sqrt{n}$ then yields the claimed $\Omega(\sqrt{n})$ adaptive one-sided lower bound for a constant $\varepsilon$ (Theorem 4.28).
1.3.3 Upper bounds
Bottleneck routing from a violation matching.
Our algorithmic upper bounds use the following combinatorial fact: distance to monotonicity is witnessed by a large matching of violated comparable pairs in the transitive closure. Concretely, if $f$ is $\varepsilon$-far, then the violation graph (on comparable pairs) contains a matching $M$ of size $|M|\geq\varepsilon n$ (via Konig’s theorem), so there are many disjoint source–sink violations $(u\!\rightsquigarrow\!v,\ f(u)>f(v))$. The difficulty is that these violations typically live in the transitive closure, denoted by $\mathrm{TC}(G)$, while the input provides only the transitive reduction (Hasse diagram), whose edge set can be much smaller and may contain very few direct violations. The key observation is that if direct violating edges are scarce, then they act as bottlenecks through which many matched violations are routed, forcing a large supply of violated pairs that can be detected by simple sampling.
The main combinatorial step is to partition the violation matching $M$ according to a first violating edge in the transitive reduction. Fix, for each $(u,v)\in M$, an arbitrary directed path $P_{uv}$ in the transitive reduction from $u$ to $v$. Since $f(u)>f(v)$, along that path there must be a first edge $(x,y)$ where the inequality flips, i.e., a violating edge of the transitive reduction. Assign $(u,v)$ to this first violating edge, and thereby partition
$M=\biguplus_{e\in F_{f}}M_{e},\qquad F_{f}:=\{(x,y)\in E:\ f(x)>f(y)\}.$
This assignment has two consequences.
  1. 1.
    It converts global violations in $\mathrm{TC}(G)$ into local witness edges in $E$.
  2. 2.
    Because $M$ is a matching, the sets of left endpoints $\{A_{e}\}$ and right endpoints $\{B_{e}\}$ of the parts $M_{e}$ are disjoint across different bottlenecks $e$. This disjointness is the main technical reason the probabilistic analysis can be carried out globally.
From here, Section 5 develops two testers (Algorithms 1 and 2), each using this dichotomy in a different way and proved in Theorems 5.1 and 5.2.
(i) A transitive-reduction/transitive-closure hybrid tester: $O\!\left(\sqrt{m\ell/(\varepsilon n)}\right)$ queries.
The first tester (Algorithm 1, see also Section 5.1) samples a moderate number of edges from the transitive reduction $E$ and, independently, a moderate number of comparable pairs from $E(\mathrm{TC}(G))$. The analysis uses the following dichotomy:
  • If many edges of the transitive reduction are violated (i.e., $|F_{f}|$ is large), then uniform sampling from $E$ finds a witness quickly.
  • If few edges of the transitive reduction are violated, then the bottleneck partition above implies a quadratic blow-up in the number of violated comparable pairs in the closure: each bottleneck edge $e$ with a large bucket $M_{e}$ certifies many violated pairs of the form $(a,b)$ with $a\in A_{e}$, $b\in B_{e}$. The median trick used in the proof formalizes this amplification: by splitting endpoints around a median threshold $t_{e}$, one obtains large subsets $A^{+}_{e}\subseteq A_{e}$, $B^{-}_{e}\subseteq B_{e}$ such that every $(a,b)\in A^{+}_{e}\times B^{-}_{e}$ is a violated pair in $\mathrm{TC}(G)$. Thus each $e$ contributes $\Omega(|M_{e}|^{2})$ violated comparable pairs, and summing (plus Cauchy–Schwarz) yields a lower bound of order $|M|^{2}/|F_{f}|$ on the number of violated closure edges.
Balancing these two regimes with an optimized threshold $\tau$ leads to a rejection probability bounded away from $1/2$ using
$q\asymp\sqrt{\frac{m\ell}{\varepsilon n}}$
samples in each loop (up to constants). This is a tradeoff between easy local witnesses and many global witnesses: if one cannot see violations locally, they must manifest globally in abundance, and vice versa. This bound can beat the classical $O(\sqrt{n/\varepsilon})$ tester on sparse reachability structures, precisely when $m\ell=o(n^{3})$ for constant $\varepsilon$.
(ii) A two-stage tester independent of $\ell$: $O(m^{1/3}/\varepsilon^{2/3})$ queries.
The second tester (Algorithm 2, see also Section 5.2) removes dependence on $\ell$ altogether by avoiding uniform sampling from $\mathrm{TC}(G)$. Instead, after a first stage that samples edges of the transitive reduction (to catch the many-violating-edges case), it runs a second stage that samples two random multisets of vertices $Q_{L},Q_{R}$ and then checks all comparable pairs between them (reachability tests and scanning $Q_{L}\times Q_{R}$ are free in the model). Here the bottleneck partition yields a stronger conclusion: the disjoint endpoint structure means that if there are few violating edges, then the matching $M$ must be split across few buckets, so some bucket $M_{e}$ must be large; and if $Q_{L}$ hits $A^{+}_{e}$ and $Q_{R}$ hits $B^{-}_{e}$ for the same $e$, the tester certifiably rejects.
The technical obstacle in making this quantitative is dependence: sampling vertices with replacement creates complicated correlations when analyzing the event “for some $e$, we hit both sides of $A^{+}_{e}\times B^{-}_{e}$.” Section 5 resolves this with a standard Poissonization argument. By coupling the multinomial sampling counts with independent Poisson variables, the analysis turns the success events for different bottlenecks $e$ into (essentially) independent events, exactly because the $A^{+}_{e}$’s are disjoint and the $B^{-}_{e}$’s are disjoint. This produces a product bound and again leverages Cauchy–Schwarz to relate $\sum_{e}|M_{e}|^{2}$ to $|M|^{2}/|F_{f}|$. Choosing the stage sizes to balance the regime in which many edges of the transitive reduction are violated with the few-bottlenecks regime yields
$q\asymp\frac{m^{1/3}}{\varepsilon^{2/3}}$
queries overall.
1.4 Related work
Monotonicity testing.
Monotonicity testing on the Boolean hypercube was initiated by Goldreich et al. [21]. For product domains, subsequent work gave improved testers, near-optimal bounds for hypercubes and hypergrids, and lower bounds over product domains and hypergrids [11, 12, 16, 5, 10, 13, 14, 15]; connections to isoperimetric inequalities for the Boolean cube were developed in [26]. For general posets, testing was initiated by Fischer et al. [19], and a complementary line used transitive-closure spanners (TC-spanners) and reachability sparsification to design testers on structured graph families [7, 28]. Beyond standard testing, distance estimation and tolerant testing were studied in [1, 27], while local reconstruction of monotone functions and its limitations appear in [6, 31]. More recently, for real-valued functions on product domains, monotonicity has been connected to isoperimetric inequalities for ordered ranges [9, 12]. In contrast, for explicit-DAG monotonicity testing with real-valued ranges, prior lower bounds in both the non-adaptive and adaptive models remained far from the $\sqrt{n}$ scale.
Graph-theoretic background.
Graphs whose edges can be partitioned into induced matchings (Ruzsa–Szemerédi graphs) have been extensively studied in extremal graph theory since [30]. In the linear regime $r=cn$, the number of induced matchings $t$ is tightly constrained in several ranges of $c$: Fox, Huang, and Sudakov show that $t=O(1)$ for $c>1/4$, that $t=\Theta(\log n)$ for $c=1/4$, and that $t=O(n/\log n)$ for every fixed $1/5+\varepsilon\leq c<1/4$ [20]. On the other hand, for every fixed $c<1/4$, Fischer et al. construct graphs with induced matchings of size $cn$ and with $t=n^{\Omega(1/\log\log n)}$ such matchings, which underlie their Boolean monotonicity-testing lower bound on general posets [19]. If one allows slightly sublinear matchings instead, Alon, Moitra, and Sudakov construct nearly complete graphs whose edges decompose into induced matchings of size $n^{1-o(1)}$, with total edge count $\binom{n}{2}-o(n^{2})$ [3].
Decompositions into positive matchings have also been studied under the name of positive matching decompositions [18]. In that line of work, the natural question is whether the edge set of a graph can be partitioned into positive matchings, and if so, into how many parts. Concretely, positive matchings are characterized by the absence of $M$-alternating closed walks (where $M$ denotes the matching under consideration) on the $M$-saturated vertices [18]; in bipartite graphs this is equivalent to being alternating-cycle-free (uniquely restricted matching) [24]. For our purposes, PMRS is the real-valued analogue of RS packings on bipartite graphs, and positivity is exactly the condition that allows a matching to appear as the set of violated edges.
1.5 Organization.
Section 2 introduces notation and basic facts. Section 3 develops the non-adaptive two-sided lower bounds and explicit PMRS constructions. Section 4 proves adaptive one-sided lower bounds. Section 5 collects the algorithmic upper bounds parameterized by $m$ and $\ell$.
2 Preliminaries
2.1 Graph notation
Let $G=(V,E)$ be a directed graph with $n:=|V|$. For disjoint vertex sets $S,T\subseteq V$, let $e(S,T)$ denote the number of edges $(u,v)\in E$ with $u\in S$ and $v\in T$. We write $u\leadsto v$ if there is a directed path from $u$ to $v$ in $G$. The transitive closure $\mathrm{TC}(G)$ has edge set
$E(\mathrm{TC}(G)):=\{(u,v)\in V\times V:u\leadsto v\},$
and we write $\ell:=|E(\mathrm{TC}(G))|$. For a DAG $G$, the transitive reduction (Hasse diagram) is the unique DAG $\mathrm{TR}(G)=(V,E_{R})$ with $\mathrm{TC}(\mathrm{TR}(G))=\mathrm{TC}(G)$ and minimal edge set. Since monotonicity depends only on reachability, we may replace $G$ by $\mathrm{TR}(G)$ without loss of generality and hence assume $G$ is transitively reduced; we write $m:=|E|$.
2.2 Positive matchings
Let $\Gamma=(V,E)$ be an undirected graph and let $M\subseteq E$ be a matching.
Definition 2.1 (Positive matching [18]).
The matching $M$ is positive (with respect to $\Gamma$) if there exists a weight function $w:V\to\mathbb{R}$ such that for every edge $e=\{u,v\}\in E$,
$w(u)+w(v)>0\quad\Longleftrightarrow\quad e\in M,$
equivalently: $w(u)+w(v)>0$ for $e\in M$ and $w(u)+w(v)\leq 0$ for $e\notin M$. [^1] Some references require $w(u)+w(v)<0$ for $e\notin M$. For finite graphs, the two formulations are equivalent: a sufficiently small perturbation of a witness $w$ makes all non-matching sums strictly negative without changing the signs on matching edges.
We recall alternating walks and the characterization of positive matchings via alternating closed walks, and its bipartite specialization to alternating-cycle-free (uniquely restricted) matchings. Let $V(M):=\{v\in V:v\text{ is incident to an edge of }M\}$.
Definition 2.2 (Alternating walks and cycles).
Let $\Gamma=(V,E)$ be a graph and $M\subseteq E$ a matching. An $M$-alternating walk is a walk whose edges alternate between $M$ and $E\setminus M$. An $M$-alternating closed walk is an alternating walk that starts and ends at the same vertex. An $M$-alternating cycle is an alternating walk that is a (simple) cycle.
Theorem 2.3 (Characterization of positive matchings [18, 24]).
A matching $M$ in $\Gamma$ is positive if and only if the induced subgraph $\Gamma[V(M)]$ contains no $M$-alternating closed walk. In bipartite graphs this is equivalent to the absence of an $M$-alternating cycle, and such matchings are also known as uniquely restricted.
2.3 Monotonicity and violations
A function $f:V\to\mathbb{R}$ is monotone (order-preserving) if $f(u)\leq f(v)$ for every $(u,v)\in E(\mathrm{TC}(G))$. Define the distance to monotonicity
$d_{\mathrm{mon}}(f):=\min_{g:V\to\mathbb{R}\ \text{monotone}}|\{v\in V:\ f(v)\neq g(v)\}|.$
For $\varepsilon>0$, we say that $f$ is $\varepsilon$-far from monotonicity if $d_{\mathrm{mon}}(f)>\varepsilon n$.
A violating pair for $f$ is an edge $(u,v)\in E(\mathrm{TC}(G))$ with $f(u)>f(v)$; when $(u,v)\in E$ we call it a violating edge. The violation graph $G_{f}$ is the bipartite graph with left part $V_{L}$ (a copy of $V$), right part $V_{R}$ (another copy of $V$), and an edge $\{u_{L},v_{R}\}$ whenever $u\leadsto v$ in $G$ and $f(u)>f(v)$. We write $\nu(H)$ for the maximum matching size of a graph $H$. By the standard vertex-cover characterization (Kőnig’s theorem), $d_{\mathrm{mon}}(f)$ equals the minimum vertex cover size of $G_{f}$ and hence $\nu(G_{f})$ [19].
2.4 Testers and query complexity
We consider randomized algorithms with oracle access to $f:V\to\mathbb{R}$. A tester queries $f$ at selected vertices and, based on the query answers (and its internal randomness), outputs accept or reject. We require completeness and soundness: for every monotone $f$, the tester accepts with probability at least $2/3$ (and with probability $1$ for one-sided testers), and for every $f$ that is $\varepsilon$-far from monotone, it rejects with probability at least $2/3$. A tester is non-adaptive if its query set depends only on its internal randomness, and one-sided if it always accepts every monotone $f$. We work in the standard massively parameterized model: the input DAG $G$ is given explicitly and only oracle queries to $f$ are counted. Hence any computation on $G$ (including computing $\mathrm{TR}(G)$, $\mathrm{TC}(G)$, $\ell$, reachability tests, and sampling uniformly from $E$ or $E(\mathrm{TC}(G))$) is free.
3 Non-Adaptive Two-Sided Lower Bounds
This section is organized as follows. We first relate positive matchings to violation patterns in bipartite DAGs and define PMRS families (Sections 3.1 and 3.2). We then prove non-adaptive two-sided lower bounds from PMRS (Section 3.3), construct explicit PMRS graphs (Section 3.4), and combine these ingredients to obtain the main near-$\sqrt{n/\varepsilon}$ lower bound (Section 3.5). Finally, we discuss how the argument changes for bounded ordered ranges (Section 3.6).
3.1 Positive matchings and violations
We recall the basic definitions and known characterizations of positive matchings in Section 2. Here we relate them to violation patterns in bipartite DAGs.
In the bipartite setting, let $U=(L,R;E)$ be an undirected bipartite graph, and let $G:=P(U)$ be the bipartite DAG obtained by orienting each edge from $L$ to $R$.
Lemma 3.1 (Violations exactly on a matching).
Let $U=(L,R;E)$ be bipartite and let $M\subseteq E$ be a matching. The following are equivalent:
  1. 1.
    $M$ is a positive matching in $U$.
  2. 2.
    There exists a function $f:L\cup R\to\mathbb{R}$ such that for every edge $(\ell,r)\in E$,
    $f(\ell)>f(r)\quad\Longleftrightarrow\quad(\ell,r)\in M.$
    In other words, the violated edges of $f$ on the bipartite DAG $G$ are exactly the matching edges $M$.
Proof.
[…]
$f(\ell)=w(\ell)\quad(\ell\in L),\qquad f(r)=-w(r)\quad(r\in R).$
Then for $(\ell,r)\in E$,
$f(\ell)-f(r)=w(\ell)+w(r),$
so $f(\ell)>f(r)$ iff $(\ell,r)\in M$.
[…]
Lemma 3.2 (Submatchings preserve positivity).
Let $\Gamma=(V,E)$ be a graph and let $M\subseteq E$ be a positive matching in $\Gamma$. Then every submatching $M^{\prime}\subseteq M$ is also a positive matching in $\Gamma$.
Proof.
Let $M^{\prime}\subseteq M$ and set $V^{\prime}:=V(M^{\prime})\subseteq V(M)$. Suppose for contradiction that $M^{\prime}$ is not positive. By Theorem 2.3, the induced subgraph $\Gamma[V^{\prime}]$ contains an $M^{\prime}$-alternating closed walk $W$. Since $M$ is a matching and $M^{\prime}\subseteq M$, no edge of $M\setminus M^{\prime}$ has an endpoint in $V^{\prime}$, and thus $M\cap E(\Gamma[V^{\prime}])=M^{\prime}$. Therefore $W$ is also an $M$-alternating closed walk contained in $\Gamma[V(M)]$, contradicting the positivity of $M$ (again by Theorem 2.3). ∎
3.2 A positive-matching analogue of RS
We now formalize PMRS families, a relaxation of RS graphs obtained by replacing induced matchings with positive matchings. This notion underlies the lower-bound constructions in the sequel.
Definition 3.3 (PMRS: Positive-Matching RS family).
Fix $\varepsilon\in(0,1/2)$ and $n_{0}\in\mathbb{N}$. A bipartite graph $U=(L,R;E)$ with $|L|=|R|=n_{0}$ is $(s,\varepsilon)$-PMRS if there exist matchings $M_{1},\dots,M_{s}\subseteq E$ such that:
  1. 1.
    (Edge-disjoint) $M_{i}\cap M_{j}=\emptyset$ for all $i\neq j$.
  2. 2.
    (Linear size) $|M_{i}|\geq\varepsilon n_{0}$ for all $i$.
  3. 3.
    (Positive) Each $M_{i}$ is a positive matching in $U$.
By Theorem 2.3, in bipartite graphs positivity is equivalent to being alternating-cycle-free (a.k.a. uniquely restricted), so PMRS can be viewed as a packing of large edge-disjoint uniquely restricted matchings.
If $U=(L,R;E)$ is $(s,\varepsilon)$-PMRS with $|L|=|R|=n_{0}$, then $|E|\geq\sum_{i}|M_{i}|\geq\varepsilon n_{0}s,$ hence $s\leq|E|/(\varepsilon n_{0})\leq n_{0}/\varepsilon$. Our explicit construction combined with refinement attains $s=n_{0}^{1-o(1)}/\varepsilon$, approaching this ceiling up to subpolynomial factors.
3.3 Non-adaptive two-sided lower bounds from PMRS
We begin by proving a non-adaptive two-sided lower bound from PMRS: for any $(s,4\varepsilon)$-PMRS graph, any non-adaptive two-sided tester requires $\Omega(\sqrt{s})$ queries. The proof builds a YES/NO distribution using a margin-normalized witness and a correlated-noise construction, which we set up below.
3.3.1 Farness from a large violated matching
Lemma 3.4 (A violated matching forces $\varepsilon$-farness).
Let $G=P(U)$ for some bipartite graph $U$, and let $f:V\to\mathbb{R}$ be a function on $G$. If the set of violated edges contains a matching $M$ of size $|M|\geq\varepsilon|V|$, then $f$ is $\varepsilon$-far from monotone.
Proof.
Each violated edge $(\ell,r)\in M$ requires changing $f(\ell)$ or $f(r)$ in any monotone correction, since the inequality $f(\ell)\leq f(r)$ depends only on these two endpoints. Because $M$ is a matching, all endpoints are distinct, so at least $|M|$ vertices must be modified. Hence the normalized distance is at least $|M|/|V|\geq\varepsilon$. ∎
3.3.2 Margin and a slack monotone base
Fix an index $i\in[s]$ and let $M_{i}$ be the corresponding matching in an $(s,4\varepsilon)$-PMRS graph $U=(L,R;E)$. Let $w_{i}:L\cup R\to\mathbb{R}$ be any weight function certifying that $M_{i}$ is positive, i.e., $w_{i}(\ell)+w_{i}(r)>0$ for $(\ell,r)\in M_{i}$ and $w_{i}(\ell)+w_{i}(r)\leq 0$ for $(\ell,r)\notin M_{i}$. Define
$\delta_{i}:=\min_{e=(\ell,r)\in M_{i}}\bigl(w_{i}(\ell)+w_{i}(r)\bigr).$
Since $M_{i}$ is finite and each term is positive, we have $\delta_{i}>0$. Define the normalized witness
$\widetilde{w}_{i}(v):=w_{i}(v)/\delta_{i}-\tfrac{1}{4}\qquad\text{for all }v\in L\cup R.$
Then for every edge $(\ell,r)\in E$,
[…]
Define $f_{i}:L\cup R\to\mathbb{R}$ by $f_{i}(\ell)=\widetilde{w}_{i}(\ell)$ for $\ell\in L$ and $f_{i}(r)=-\widetilde{w}_{i}(r)$ for $r\in R$. Then, for every $(\ell,r)\in E$,
$f_{i}(\ell)-f_{i}(r)=\widetilde{w}_{i}(\ell)+\widetilde{w}_{i}(r),$
so by (1) we have $f_{i}(\ell)-f_{i}(r)\geq+\tfrac{1}{2}$ on $M_{i}$ and $\leq-\tfrac{1}{2}$ on $E\setminus M_{i}$.
Let $L_{i}\subseteq L$ and $R_{i}\subseteq R$ be the sets of endpoints of $M_{i}$. For each matching edge $e=(\ell,r)\in M_{i}$, define $\Delta_{e}:=f_{i}(\ell)-f_{i}(r)$, so by (1) we have $\Delta_{e}\geq\tfrac{1}{2}$. We convert $f_{i}$ into a monotone function with slack by shifting matching endpoints:
[…]
Here, for $\ell\in L_{i}$ (resp. $r\in R_{i}$), $e$ denotes the unique edge of $M_{i}$ incident to $\ell$ (resp. $r$); this is well-defined since $M_{i}$ is a matching. Then for every edge $(\ell,r)\in E$ we have
$(\ell,r)\in M_{i}\implies g_{i}(\ell)=g_{i}(r),\qquad(\ell,r)\notin M_{i}\implies g_{i}(\ell)\leq g_{i}(r)-\tfrac{1}{2}.$ (3)
Indeed, for non-matching edges we have $f_{i}(\ell)-f_{i}(r)\leq-\tfrac{1}{2}$ by (1), and the shifts in (2) only decrease the left endpoint (if any) and increase the right endpoint (if any), so the difference can only decrease. In particular, $g_{i}$ is monotone on the bipartite DAG $G$, and every non-matching edge has slack at least $1/2$.
3.3.3 YES/NO distributions
We now define two distributions over functions. Fix the noise alphabet size $K:=4$ and step size $h:=1/8$. For each matching edge $e=(\ell,r)\in M_{i}$, sample $Z_{e}\sim\mathrm{Unif}(\{0,1,2,3\})$ independently, and define the noise value $\eta_{e}:=hZ_{e}\in\{0,h,2h,3h\}$.
YES distribution $\mathcal{D}_{+}$. Sample $i\sim\mathrm{Unif}([s])$, build $g_{i}$ as in (2), sample $\{Z_{e}\}_{e\in M_{i}}$, and define $F_{i}^{+}:L\cup R\to\mathbb{R}$ by
$F_{i}^{+}(v)=g_{i}(v)\quad\text{for }v\notin L_{i}\cup R_{i},\qquad F_{i}^{+}(\ell)=g_{i}(\ell)+\eta_{e},\quad F_{i}^{+}(r)=g_{i}(r)+\eta_{e}\quad\text{for }e=(\ell,r)\in M_{i}.$
NO distribution $\mathcal{D}_{-}$. Sample $i$ and $\{Z_{e}\}_{e\in M_{i}}$ as above, and define $F_{i}^{-}$ by
$F_{i}^{-}(v)=g_{i}(v)\quad\text{for }v\notin L_{i}\cup R_{i},\qquad F_{i}^{-}(\ell)=g_{i}(\ell)+\eta_{e},\quad F_{i}^{-}(r)=g_{i}(r)+\eta^{\prime}_{e}\quad\text{for }e=(\ell,r)\in M_{i},$
where $\eta^{\prime}_{e}:=h\cdot((Z_{e}-1)\bmod 4)$ (i.e., we cyclically shift the noise on the right endpoint).
Note that in both distributions, the marginal noise on a single endpoint is the same: it is uniform over $\{0,h,2h,3h\}$. The only difference is the correlation between the two endpoints of the same matching edge.
3.3.4 Completeness, soundness, and indistinguishability
Lemma 3.5 (Completeness).
Every function in the support of $\mathcal{D}_{+}$ is monotone on $G$.
Proof.
Fix $i$ and consider any edge $(\ell,r)\in E$.
If $(\ell,r)\in M_{i}$, then $g_{i}(\ell)=g_{i}(r)$ by (3) and the same noise $\eta_{e}$ is added to both endpoints, so $F_{i}^{+}(\ell)=F_{i}^{+}(r)$.
If $(\ell,r)\notin M_{i}$, then $g_{i}(\ell)\leq g_{i}(r)-1/2$ by (3). The noise values lie in $[0,3h]$, hence
$F_{i}^{+}(\ell)-F_{i}^{+}(r)\leq(g_{i}(\ell)-g_{i}(r))+3h\leq-\tfrac{1}{2}+\tfrac{3}{8}=-\tfrac{1}{8}<0.$
Thus $F_{i}^{+}(\ell)\leq F_{i}^{+}(r)$ for every edge, i.e., $F_{i}^{+}$ is monotone. ∎
Lemma 3.6 (Many violations in the NO distribution).
Fix $i$ and draw $F_{i}^{-}\sim\mathcal{D}_{-}$. Then an edge $e=(\ell,r)\in M_{i}$ is violated by $F_{i}^{-}$ iff $Z_{e}\neq 0$. Consequently, the violated edges contain a matching of size
$X_{i}:=|\{e\in M_{i}:Z_{e}\neq 0\}|\sim\mathrm{Bin}(|M_{i}|,3/4).$
In particular, if $X_{i}\geq 2\varepsilon n_{0}$, then $F_{i}^{-}$ is $\varepsilon$-far from monotone (by Section 3.3.1). Moreover, if $|M_{i}|\geq 4\varepsilon n_{0}$ (e.g., when $U$ is $(s,4\varepsilon)$-PMRS), then for all sufficiently large $n_{0}$,
$\Pr[X_{i}\geq 2\varepsilon n_{0}]\ \geq\ 1-\exp(-\Omega(\varepsilon n_{0}))\ \geq\ 9/10.$
Proof.
For $e=(\ell,r)\in M_{i}$, we have $g_{i}(\ell)=g_{i}(r)$, hence
$F_{i}^{-}(\ell)-F_{i}^{-}(r)=\eta_{e}-\eta^{\prime}_{e}.$
If $Z_{e}\in\{1,2,3\}$, then $\eta_{e}-\eta^{\prime}_{e}=h>0$ so $e$ is violated. If $Z_{e}=0$, then $\eta_{e}-\eta^{\prime}_{e}=0-3h<0$ so $e$ is not violated. Thus the violated edges are exactly $\{e\in M_{i}:Z_{e}\neq 0\}$, which is a submatching of $M_{i}$ of size $X_{i}$.
If $X_{i}\geq 2\varepsilon n_{0}$, then $X_{i}\geq\varepsilon|V|$ because $|V|=2n_{0}$, and Section 3.3.1 implies $\varepsilon$-farness.
Finally, if $|M_{i}|\geq 4\varepsilon n_{0}$, then $\mu:=\mathbb{E}[X_{i}]=(3/4)|M_{i}|\geq 3\varepsilon n_{0}$. A Chernoff bound gives
$\Pr[X_{i}<2\varepsilon n_{0}]=\Pr\!\left[X_{i}<\left(1-\tfrac{1}{3}\right)\mu\right]\leq\exp\!\left(-\frac{(1/3)^{2}\mu}{2}\right)\leq\exp\!\left(-\frac{\varepsilon n_{0}}{6}\right),$
which is at most $1/10$ for all sufficiently large $n_{0}$. ∎
Lemma 3.7 (No queried matching edge $\Rightarrow$ identical transcripts).
Fix $i\in[s]$ and a non-adaptive query set $Q\subseteq L\cup R$. If $Q$ does not contain both endpoints of any edge in $M_{i}$, then the joint distribution of the oracle answers on $Q$ is identical under $F_{i}^{+}$ and $F_{i}^{-}$.
Proof.
Under the assumption, each matching edge contributes to the transcript through at most one queried endpoint. For any queried endpoint of a matching edge $e$, the observed noise is a uniform element of $\{0,h,2h,3h\}$ both in $F_{i}^{+}$ and in $F_{i}^{-}$ (because $Z_{e}$ is uniform and the map $z\mapsto(z-1)\bmod 4$ is a permutation). Across distinct edges, the noises are independent in both distributions. Hence the induced joint distribution of all answers on $Q$ is identical. ∎
Theorem 3.8 (Non-adaptive two-sided $\Omega(\sqrt{s})$ lower bound).
Let $\varepsilon\in(0,1/8]$ and let $U=(L,R;E)$ be any $(s,4\varepsilon)$-PMRS graph with $|L|=|R|=n_{0}$. Set $G:=P(U)$. Then any non-adaptive two-sided $\varepsilon$-tester for monotonicity on the bipartite DAG $G$ with range $\mathbb{R}$ must make $q=\Omega(\sqrt{s})$ queries.
Proof.
By Yao’s minimax principle, it suffices to consider deterministic non-adaptive testers. Fix such an algorithm $\mathcal{A}$, and let $Q$ be its query set with $|Q|=q$. Define
$E(Q):=E\cap(Q\cap L)\times(Q\cap R),$
the set of edges whose both endpoints are queried.
Condition on the choice of $i\in[s]$. If $E(Q)\cap M_{i}=\emptyset$, then $Q$ contains no full matching edge from $M_{i}$, so by Section 3.3.4 the transcript distribution is identical under $F_{i}^{+}$ and $F_{i}^{-}$. Therefore, for this fixed $Q$,
$\mathrm{TV}\big(\mathrm{Tr}(\mathcal{D}_{+},Q),\mathrm{Tr}(\mathcal{D}_{-},Q)\big)\ \leq\ \Pr_{i\sim[s]}\big[E(Q)\cap M_{i}\neq\emptyset\big].$
Since the matchings $M_{1},\dots,M_{s}$ are edge-disjoint, each edge in $E(Q)$ belongs to at most one matching. Hence the above event can occur for at most $|E(Q)|$ indices $i$, implying
$\mathrm{TV}\big(\mathrm{Tr}(\mathcal{D}_{+},Q),\mathrm{Tr}(\mathcal{D}_{-},Q)\big)\ \leq\ \frac{|E(Q)|}{s}\ \leq\ \frac{|Q\cap L|\cdot|Q\cap R|}{s}\ \leq\ \frac{q^{2}}{4s}.$ (4)
Now, $\mathcal{D}_{+}$ is supported on monotone functions (Section 3.3.4), so completeness implies
$\Pr[\mathcal{A}\text{ accepts }f\sim\mathcal{D}_{+}]\geq 2/3.$
On the other hand, by Section 3.3.4, with probability at least $9/10$ (over $f\sim\mathcal{D}_{-}$) the sampled function is $\varepsilon$-far; thus soundness implies
$\Pr[\mathcal{A}\text{ rejects }f\sim\mathcal{D}_{-}]\ \geq\ (9/10)\cdot(2/3)\ =\ 3/5.$
Equivalently, $\Pr[\mathcal{A}\text{ accepts }f\sim\mathcal{D}_{-}]\leq 2/5$.
Therefore, $\mathcal{A}$ distinguishes $\mathcal{D}_{+}$ from $\mathcal{D}_{-}$ with constant bias:
$\Pr[\mathcal{A}\text{ accepts }\mathcal{D}_{+}]-\Pr[\mathcal{A}\text{ accepts }\mathcal{D}_{-}]\ \geq\ \frac{2}{3}-\frac{2}{5}\ =\ \frac{4}{15}.$
Any such bias is upper bounded by the total variation distance in (4), hence $q^{2}/(4s)\geq 4/15$, which yields $q=\Omega(\sqrt{s})$. ∎
3.4 An explicit PMRS construction
Fix an integer $k\geq 2$ and a parameter $N\in\mathbb{N}$. Write $N]=\{0,1,\dots,N-1\}$ and denote vectors in $N]^{k}$ by bold letters. Let $\|\mathbf{a}\|^{2}=\sum_{i=1}^{k}a_{i}^{2}$ and $\mathbf{a}\cdot\mathbf{x}=\sum_{i=1}^{k}a_{i}x_{i}$.
3.4.1 Vertex sets
Let $L$ and $R$ be two disjoint copies of $N]^{k}\times[N^{2}]$:
$L=\{(\mathbf{x},z)_{L}:\mathbf{x}\in[N]^{k},\ z\in[N^{2}]\},\qquad R=\{(\mathbf{y},t)_{R}:\mathbf{y}\in[N]^{k},\ t\in[N^{2}]\}.$
Then
$n_{0}:=|L|=|R|=N^{k}\cdot N^{2}=N^{k+2}.$
3.4.2 A family of shift vectors
Fix a constant $\alpha\in(0,1)$ (to be chosen as a function of $\varepsilon$ and $k$) and set
$P:=\lfloor\alpha N\rfloor,\qquad A:=\{0,1,\dots,P\}^{k}\setminus\{\mathbf{0}\},\qquad s:=|A|=(P+1)^{k}-1.$
3.4.3 Matchings and the underlying bipartite graph
For each $\mathbf{a}\in A$, define a set of edges $M_{\mathbf{a}}\subseteq L\times R$ by
$M_{\mathbf{a}}:=\Bigl\{\bigl((\mathbf{x},z)_{L},\ (\mathbf{x}+\mathbf{a},\ z+\|\mathbf{a}\|^{2})_{R}\bigr)\;:\;\mathbf{x}\in[N]^{k},\ z\in[N^{2}],\ \mathbf{x}+\mathbf{a}\in[N]^{k},\ z+\|\mathbf{a}\|^{2}\in[N^{2}]\Bigr\}.$
Let
$E:=\bigcup_{\mathbf{a}\in A}M_{\mathbf{a}}\quad\text{and}\quad U:=(L,R;E).$
Theorem 3.9 (Polynomial-size PMRS).
Fix $\varepsilon\in(0,1/2)$ and an integer $k\geq 2$. There exists a choice of $\alpha=\alpha(\varepsilon,k)>0$ such that for all sufficiently large $N$ (so that $P\geq 1$), the bipartite graph $U$ above is an $(s,\varepsilon)$-PMRS with $n_{0}=N^{k+2}$ and
$s=\Omega\!\bigl(n_{0}^{k/(k+2)}\bigr).$
In particular, taking $k=2$ yields $s=\Omega(\sqrt{n_{0}})$.
Proof.
We verify the three PMRS conditions.
(1) Each $M_{\mathbf{a}}$ is a matching. Fix $\mathbf{a}\in A$. Each left vertex $(\mathbf{x},z)_{L}$ is incident to at most one edge of $M_{\mathbf{a}}$, namely to $(\mathbf{x}+\mathbf{a},z+\|\mathbf{a}\|^{2})_{R}$ if that vertex lies in $R$. Conversely, each right vertex has at most one preimage under this translation map. Hence $M_{\mathbf{a}}$ is a matching.
(2) Edge-disjointness. Every edge $e\in E$ belongs to a unique $M_{\mathbf{b}}$: indeed, if
$e=\bigl((\mathbf{x},z)_{L},(\mathbf{y},t)_{R}\bigr)\in E,$
then necessarily $\mathbf{y}-\mathbf{x}=\mathbf{b}$ for some $\mathbf{b}\in A$, and this $\mathbf{b}$ is uniquely determined by the endpoints. Thus the matchings $\{M_{\mathbf{a}}\}_{\mathbf{a}\in A}$ are pairwise edge-disjoint.
(3) Linear size. For $\mathbf{a}\in A$,
$|M_{\mathbf{a}}|=\Bigl(\prod_{i=1}^{k}(N-a_{i})\Bigr)\,(N^{2}-\|\mathbf{a}\|^{2}).$
Since $0\leq a_{i}\leq P\leq\alpha N$ and $\|\mathbf{a}\|^{2}\leq kP^{2}\leq k\alpha^{2}N^{2}$, we have
$|M_{\mathbf{a}}|\geq((1-\alpha)N)^{k}\cdot(1-k\alpha^{2})N^{2}=n_{0}\,(1-\alpha)^{k}(1-k\alpha^{2}).$
Choose $\alpha=\alpha(\varepsilon,k)>0$ small enough so that
$(1-\alpha)^{k}(1-k\alpha^{2})\ \geq\ \varepsilon,$
which is possible because the left-hand side tends to $1$ as $\alpha\to 0$. Then $|M_{\mathbf{a}}|\geq\varepsilon n_{0}$ for all $\mathbf{a}\in A$.
(4) Positivity of each $M_{\mathbf{a}}$. Fix $\mathbf{a}\in A$. Define a weight function $w_{\mathbf{a}}:L\cup R\to\mathbb{R}$ by
$w_{\mathbf{a}}\bigl((\mathbf{x},z)_{L}\bigr):=z-2\,\mathbf{a}\cdot\mathbf{x},\qquad w_{\mathbf{a}}\bigl((\mathbf{y},t)_{R}\bigr):=-t+2\,\mathbf{a}\cdot\mathbf{y}-\|\mathbf{a}\|^{2}+\tfrac{1}{2}.$
Consider any edge $e\in E$. By construction, $e$ lies in a unique $M_{\mathbf{b}}$ for some $\mathbf{b}\in A$, and has the form
$e=\bigl((\mathbf{x},z)_{L},\ (\mathbf{x}+\mathbf{b},z+\|\mathbf{b}\|^{2})_{R}\bigr).$
A direct calculation gives
$w_{\mathbf{a}}\bigl((\mathbf{x},z)_{L}\bigr)+w_{\mathbf{a}}\bigl((\mathbf{x}+\mathbf{b},z+\|\mathbf{b}\|^{2})_{R}\bigr)$ $=\bigl(z-2\mathbf{a}\cdot\mathbf{x}\bigr)+\bigl(-(z+\|\mathbf{b}\|^{2})+2\mathbf{a}\cdot(\mathbf{x}+\mathbf{b})-\|\mathbf{a}\|^{2}+\tfrac{1}{2}\bigr)$
$=\tfrac{1}{2}-\|\mathbf{b}-\mathbf{a}\|^{2}.$
Since $\mathbf{a},\mathbf{b}$ have integer coordinates, $\|\mathbf{b}-\mathbf{a}\|^{2}=0$ iff $\mathbf{b}=\mathbf{a}$, and otherwise $\|\mathbf{b}-\mathbf{a}\|^{2}\geq 1$. Hence
$w_{\mathbf{a}}(u)+w_{\mathbf{a}}(v)>0\iff\mathbf{b}=\mathbf{a}\iff e\in M_{\mathbf{a}},$
and for every edge $e\notin M_{\mathbf{a}}$ we have $w_{\mathbf{a}}(u)+w_{\mathbf{a}}(v)\leq-\tfrac{1}{2}\leq 0$. This is exactly the definition of $M_{\mathbf{a}}$ being a positive matching in $U$.
(5) Counting matchings. Finally,
$s=|A|=(P+1)^{k}-1=\Theta(N^{k}).$
Since $n_{0}=N^{k+2}$, this yields
$s=\Theta\!\bigl(n_{0}^{k/(k+2)}\bigr),$
as claimed. ∎
3.5 Main lower bound
We now combine the PMRS lower-bound argument with the explicit construction. As a first step, we state a refinement lemma that converts larger positive matchings into many smaller ones while preserving positivity.
Lemma 3.10 (Refining a PMRS family to smaller $\varepsilon$).
Let $\varepsilon_{0}\in(0,1/2)$ and let $U=(L,R;E)$ be an $(s,\varepsilon_{0})$-PMRS graph with $|L|=|R|=n_{0}$. Then for every $\varepsilon\in(0,\varepsilon_{0}]$, the same graph contains an $(s^{\prime},\varepsilon)$-PMRS family with
$s^{\prime}\;\geq\;\sum_{i=1}^{s}\Big\lfloor\frac{|M_{i}|}{\varepsilon n_{0}}\Big\rfloor\;\geq\;s\Big\lfloor\frac{\varepsilon_{0}}{\varepsilon}\Big\rfloor.$
Proof.
Let $M_{1},\dots,M_{s}$ be the edge-disjoint positive matchings from Section 3.2. Fix $i$ and define
$t_{i}:=\big\lfloor|M_{i}|/(\varepsilon n_{0})\big\rfloor.$
Partition $M_{i}$ arbitrarily into $t_{i}$ pairwise edge-disjoint submatchings $M_{i,1},\dots,M_{i,t_{i}}$, each of size exactly $\varepsilon n_{0}$ (discarding a remainder of size $\varepsilon n_{0}$). Each $M_{i,j}$ is a matching, and it is positive by Section 3.1. Since the original $M_{i}$’s are edge-disjoint across different $i$, all $M_{i,j}$ are edge-disjoint as well. Moreover, $|M_{i}|\geq\varepsilon_{0}n_{0}$ implies $t_{i}\geq\lfloor\varepsilon_{0}/\varepsilon\rfloor$. Thus the collection $\{M_{i,j}\}$ forms an $(s^{\prime},\varepsilon)$-PMRS family with $s^{\prime}=\sum_{i}t_{i}$ and the stated lower bound on $s^{\prime}$. ∎
Corollary 3.11 ($\varepsilon$-dependent two-sided PMRS lower bound).
Let $\varepsilon_{0}\in(0,1/2)$ and let $U$ be an $(s,\varepsilon_{0})$-PMRS graph with $|L|=|R|=n_{0}$. Set $G:=P(U)$. Then for every $\varepsilon\in(0,\varepsilon_{0}/4]$, every non-adaptive two-sided $\varepsilon$-tester for monotonicity on $G$ requires
$q\;=\;\Omega\!\left(\sqrt{\frac{s\varepsilon_{0}}{\varepsilon}}\right)$
queries.
Proof.
By Section 3.5 with parameter $4\varepsilon$, $U$ contains an $(s^{\prime},4\varepsilon)$-PMRS family with $s^{\prime}\geq s\lfloor\varepsilon_{0}/(4\varepsilon)\rfloor$. Applying Theorem 3.8 to this family yields $q=\Omega(\sqrt{s^{\prime}})$, which gives the claimed bound (absorbing the floor into constants). ∎
Theorem 3.12 (Near-$\sqrt{n/\varepsilon}$ two-sided lower bounds).
Fix $\varepsilon\in(0,1/8]$. For every $\delta>0$, there exist infinitely many $n$ and bipartite DAGs on $n$ vertices for which any non-adaptive two-sided $\varepsilon$-tester needs
$q\;=\;\Omega\!\left(\frac{n^{1/2-\delta}}{\sqrt{\varepsilon}}\right)$
queries.
Proof.
Let $\varepsilon_{0}:=1/2$. Choose an integer $k\geq 2$ so that $1/(k+2)\leq\delta$. By Theorem 3.9, for all sufficiently large $N$ there exists an $(s_{0},\varepsilon_{0})$-PMRS graph with $n_{0}=N^{k+2}$ and $s_{0}=\Omega\!\bigl(n_{0}^{k/(k+2)}\bigr)$. Applying Section 3.5 yields
$q=\Omega\!\left(\sqrt{\frac{s_{0}\varepsilon_{0}}{\varepsilon}}\right)\;=\;\Omega\!\left(\frac{n_{0}^{k/(2k+4)}}{\sqrt{\varepsilon}}\right)\;=\;\Omega\!\left(\frac{n_{0}^{1/2-1/(k+2)}}{\sqrt{\varepsilon}}\right).$
Now let $n:=|V|=2n_{0}$. Since $n$ and $n_{0}$ differ only by a factor of $2$, this is
$q=\Omega\!\left(\frac{n^{1/2-1/(k+2)}}{\sqrt{\varepsilon}}\right)=\Omega\!\left(\frac{n^{1/2-\delta}}{\sqrt{\varepsilon}}\right).$
Because this construction exists for all sufficiently large $N$, it gives infinitely many values of $n$. ∎
3.6 Bounded ordered ranges and polylogarithmic ranges
So far we have worked over an unbounded totally ordered range (e.g. $\mathbb{R}$) for notational convenience, and noted that the constructions can be scaled to $\mathbb{Z}$ (and any unbounded totally ordered range). In this subsection we briefly discuss what happens if the range is a finite totally ordered set, say $B]=\{0,1,\dots,B-1\}$, and in particular when $B=\mathrm{polylog}(n)$.
The PMRS lower-bound argument does not inherently require unboundedness (two-sided).
The non-adaptive two-sided lower bound derived from an $(s,\varepsilon_{0})$-PMRS decomposition (Theorems 3.8 and 3.5) is information-theoretic: it constructs YES/NO distributions by adding a constant-alphabet noise to a slack monotone base and argues that a non-adaptive query set cannot distinguish the two distributions unless it queries both endpoints of some matching edge. The proof uses only the relative order of function values and a constant slack/margin separation; it does not rely on the range being unbounded beyond the ability to realize the prescribed pattern of inequalities. Consequently, the same argument extends to $[B]$ provided that the construction can be implemented with values lying in $[B]$ (after an order-preserving relabeling and an affine scaling/translation if needed).
Formally, it suffices that the set of values used by the YES/NO distributions admits an order-preserving injection into $[B]$. Since the construction in Section 3.3 uses a constant slack (at least $1/2$) and a constant noise alphabet (of size $4$), one may first scale to integers and then shift so that all used values are contained in $\{0,1,\dots,B-1\}$, as long as the dynamic range required by the monotone base is at most $B$.
Dynamic-range requirement of the explicit PMRS construction.
In the explicit PMRS graph from Section 3.4 (Theorem 3.9), the vertex sets are $L,R\cong[N]^{k}\times[N^{2}]$ and thus $n_{0}:=|L|=|R|=N^{k+2}$ (so the total number of vertices is $n=2n_{0}$). The positivity witness for a fixed matching index $\mathbf{a}$ is
$w_{\mathbf{a}}((\mathbf{x},z)_{L})=z-2\mathbf{a}\cdot\mathbf{x},\qquad w_{\mathbf{a}}((\mathbf{y},t)_{R})=-t+2\mathbf{a}\cdot\mathbf{y}-\|\mathbf{a}\|_{2}^{2}+\tfrac{1}{2}.$
Because $z,t\in[N^{2}]$, the witness (and hence the derived slack monotone base in Section 3.3.2) necessarily varies on a scale $\Theta(N^{2})$ (up to factors depending mildly on $k$ and the choice of $\alpha$). The subsequent normalization and constant-alphabet noise addition preserve the required comparisons while changing the value span only by constant factors. Therefore, to realize the same hard instance over a bounded range $[B]$, it is natural to require
$B\;\gtrsim\;c\,N^{2}$ (5)
for a suitable absolute constant $c>0$ (after scaling to integer values and shifting).
Specialization to $B=(\log n)^{C}$ and the resulting lower bound.
Assume $B=(\log n)^{C}$ for a fixed constant $C>0$, where $n$ denotes the total number of vertices. Since $n_{0}=\Theta(n)$, enforcing (5) by taking
$N\asymp\sqrt{B}=(\log n)^{C/2}$
forces the parameter $k$ (the “dimension” in Theorem 3.9) to grow with $n$ via $n_{0}=N^{k+2}$, namely
$k=\log_{N}n_{0}-2=\frac{\log n_{0}}{\log N}-2=\frac{2}{C}\cdot\frac{\log n}{\log\log n}-2\;=\;\Theta\!\Big(\frac{\log n}{\log\log n}\Big).$
To keep each matching $M_{\mathbf{a}}$ of linear size (Condition (2) in Theorem 3.9), the parameter $\alpha$ must shrink with $k$; a natural choice is $\alpha=\Theta(1/k)$ so that $(1-\alpha)^{k}$ remains a constant (and $1-k\alpha^{2}=1-o(1)$). With $\alpha=\Theta(1/k)$, we have $P=\lfloor\alpha N\rfloor=\Theta(N/k)$ and hence the number of matchings is
$s=(P+1)^{k}-1\approx(N/k)^{k}.$
Plugging $N=(\log n)^{C/2}$ and $k=\Theta(\log n/\log\log n)$ into $\log s\approx k\log(N/k)$ gives the asymptotic exponent
$\log s=\Big(1-\frac{2}{C}-o(1)\Big)\log n,\qquad\text{i.e.}\qquad s=n^{\,1-\frac{2}{C}-o(1)}\quad\text{(for $C>2$)}.$
(When $C=2$, one obtains $s=n^{o(1)}$; when $C<2$, $N/k\to 0$ and the above parameter choice does not yield a nontrivial PMRS packing.)
Finally, applying the $\varepsilon$-dependent PMRS lower bound (Section 3.5, with constant $\varepsilon_{0}$) to this $s$ yields, for every fixed $\varepsilon\leq\varepsilon_{0}/4$,
$q\;=\;\Omega\!\left(\sqrt{\frac{s}{\varepsilon}}\right)\;=\;\Omega\!\left(\frac{n^{\,\frac{1}{2}-\frac{1}{C}-o(1)}}{\sqrt{\varepsilon}}\right)\qquad\text{for }B=(\log n)^{C}\text{ with }C>2.$
Thus, under the explicit construction of Section 3.4, a polylogarithmic bounded range of size $(\log n)^{C}$ yields a polynomial non-adaptive two-sided lower bound precisely once $C>2$, with exponent $\tfrac{1}{2}-\tfrac{1}{C}$ (up to $o(1)$ terms).
4 Adaptive One-Sided Lower Bounds
In this section, we establish adaptive one-sided lower bounds in the explicit-DAG model, culminating in Theorem 4.28. We begin by reducing to the bipartite DAG setting.
Let $U=(L,R;E)$ be bipartite, let $n:=|L\cup R|$, and write $n_{0}:=|L|=|R|=n/2$. Let $G=P(U)$ be the DAG obtained by orienting all edges from $L$ to $R$. Then $E(\mathrm{TC}(G))=E$, so a witnessed violation is a queried edge $(\ell,r)\in E$ with $f(\ell)>f(r)$. We first record the one-sided witness principle:
Lemma 4.1 (One-sided witness principle).
Let $Q\subseteq L\cup R$ and answers $\{a_{v}\}_{v\in Q}$ satisfy $a_{\ell}\leq a_{r}$ for every queried edge $(\ell,r)\in E$ with both endpoints in $Q$. Then there exists a global monotone extension $g:L\cup R\to\mathbb{R}$ with $g|_{Q}=a$. Hence one-sided testers reject only after witnessing a violated queried edge.
Proof.
Let $A=Q\cap L$, $B=Q\cap R$, choose $a_{-}<\min(\{a_{r}:r\in B\}\cup\{0\})$ and $a_{+}>\max(\{a_{\ell}:\ell\in A\}\cup\{0\})$. Set $g(v)=a_{v}$ on $Q$, $g(\ell)=a_{-}$ on $L\setminus Q$, and $g(r)=a_{+}$ on $R\setminus Q$. Then $g(\ell)\leq g(r)$ on every edge. ∎
The rest of this section is organized as follows. We first set up the PMRS base graph and truncated Gibbs distribution (Section 4.1), prove edge-tail bounds and parameter admissibility (Section 4.2), establish farness and control of spurious queried violations (Section 4.3), and bound adaptive information leakage via a Green-function comparison (Section 4.4). Next, we combine these ingredients with decoding and Fano’s inequality (Section 4.5) to obtain a fixed-budget lower bound (Section 4.6). Finally, we instantiate parameters on an explicit dense base graph to derive the $\Omega(\sqrt{n})$ adaptive one-sided lower bound (Section 4.7).
Notation map for the adaptive proof.
The next three tables collect symbols that are reused across multiple subsections. They are not exhaustive, but cover the notation that is most frequently referenced.
(A) Global parameters and base-graph quantities.
Symbol First defined Meaning / role Mainly used in
$n,n_{0}$ opening of this section total number of vertices and side size in the bipartite base graph ($n=2n_{0}$) throughout this section
$s,\{M_{i}\}_{i=1}^{s}$ opening of Section 4.1 PMRS matching family; hidden index $I$ selects a matching Sections 4.3 and 4.5
$\delta,\Delta,m$ opening definitions in Section 4.1 min/max degrees and edge count of the base graph Sections 4.2 and 4.7
$Q$ opening definitions in Section 4.1 query-budget scale for leakage and fixed-budget lower bounds Sections 4.4 and 4.6
$\alpha,\lambda,\beta,\gamma,B$ parameter block in Section 4.1 Gibbs/hinge and truncation parameters (curvature, slope, margin, box size) all analytic subsections
$m_{\min}$ one-site curvature setup in Section 4.1 uniform one-site strong-convexity lower bound, $m_{\min}=\alpha+2\lambda\delta$ Sections 4.2 and 4.4
$m_{\mathrm{edge}}$ (19) one-dimensional edge-potential curvature lower bound Sections 4.2 and 4.6
$u$ drift-control setup in Section 4.2 tolerance level for one-sided drift/mean control events Sections 4.2 and 4.7
(B) Potentials and derived random quantities.
Symbol First defined Meaning / role Mainly used in
$\mu_{i},\ H_{i}$ opening definitions in Section 4.1 truncated Gibbs law for hidden index $i$, with Hamiltonian $H_{i}$ Section 4.1 and all downstream bounds
$W_{e,t}(d)$ (21) one-dimensional potential on difference coordinate $d=f(\ell)-f(r)$ for edge-tail analysis Section 4.2
$U_{e,t},\Psi_{e,i},R_{e,t}$ (21) decomposition of $W_{e,t}$ into quadratic part, designated hinge term, and external drift Section 4.2 and Section 4.7
$R^{\prime}_{e,t}(d)$ drift paragraph in Section 4.2 derivative of external drift along $d$ (controls edge-tail sign tests) Section 4.2 and Section 4.7
$M_{e,t}(d_{0})$ Section 4.7 conditional mean drift functional, $M_{e,t}(d_{0})=\operatorname{E}[R^{\prime}_{e,t}(d_{0})\mid\cdots]$ Section 4.7
$V(x_{\ell},x_{r})$ mean-control proof in Section 4.7 effective endpoint potential after integrating out neighbor spins on incident edges Section 4.7
$\eta_{\mathrm{strip}}$ (24) boundary-strip leakage scale controlling hard-truncation boundary effects Section 4.7
(C) Events, failure probabilities, and leakage terms.
Symbol First defined Meaning / role Mainly used in
$\mathsf{Open}_{e,t}$ mean/drift event setup in Section 4.2 edge is not fully revealed before round $t$ (at most one endpoint known) Section 4.2 and Section 4.7
$\mathsf{MeanBad}^{\pm}_{e,t}$ (17) one-sided failures of conditional mean drift control at $d_{0}=\pm\gamma/2$ Section 4.2
$p_{\mathrm{drift}}(u)$, $p_{\mathrm{mean}}(u)$ (12), (18) uniform upper bounds for drift fluctuation failure and mean-control failure Section 4.2 and Section 4.7
$\mathsf{SpurQry}$, $\mathsf{BdryQry}$ Section 4.3 events of spurious queried violations and boundary-touching queried edges Sections 4.3 and 4.6
$\eta_{\partial},\eta_{\partial,\mathcal{A}}$ Section 4.5 and Section 4.6 probabilities of $\mathsf{BdryQry}$ (global and conditioned on tester randomness) Section 4.6
$\mathbf{M}_{\tau},\mathbf{G}_{\tau}$ Section 4.4 conditioned curvature matrix and Green matrix inverse Section 4.4
$b_{u}^{(i,j)}$ Section 4.4 one-site KL influence term at site $u$ between indices $i,j$ Section 4.4
$\kappa_{Q}$ Section 4.4 per-query information leakage scale in the mutual-information bound Section 4.4 and Section 4.5
$P_{e}$ Section 4.5 decoding error probability for recovering hidden index from transcript Section 4.5 and Section 4.6
4.1 PMRS base graph and truncated Gibbs distribution
In this subsection, we define the truncated Gibbs distribution on a PMRS base graph and record the basic one-site curvature setup used later.
Assume edge-disjoint matchings $M_{1},\dots,M_{s}\subseteq E$ with
$|M_{i}|\geq\varepsilon_{0}n_{0}\qquad(\forall i\in[s]),$
for a fixed constant $\varepsilon_{0}\in(0,1/4]$. Let $I\sim\mathrm{Unif}([s])$ and define
$\delta:=\min_{v\in L\cup R}\deg(v),\qquad\Delta:=\max_{v\in L\cup R}\deg(v),\qquad m:=|E|.$
Fix an external query budget $Q\geq 1$. Set parameters $(\alpha,\lambda,\beta,\gamma,B)$ (possibly depending on $n,Q$), where:
$\alpha=\Theta(1),\qquad\lambda>0,\qquad\beta\geq 10,$
and $\gamma=\gamma(Q)$, $B=B(Q,n)$ are chosen later. Crucially, $\alpha$ is not scaled with $\Delta$.
Let
$\mathcal{K}_{B}:=[-B,B]^{L\cup R}.$
For each $i\in[s]$, define
$\mu_{i}\equiv\mu_{i}^{(Q)}(df):=\frac{1}{Z_{i}}\exp(-H_{i}(f))\,\mathbf{1}_{\mathcal{K}_{B}}(f)\,df,$
with energy
$H_{i}(f):=\$$\frac{\alpha}{2}\sum_{v\in L\cup R}f(v)^{2}+\lambda\sum_{(\ell,r)\in E}(f(\ell)-f(r))^{2}$
$+\beta\sum_{(\ell,r)\in E\setminus M_{i}}(f(\ell)-f(r)+\gamma)_{+}^{2}+\beta\sum_{(\ell,r)\in M_{i}}(\gamma-(f(\ell)-f(r)))_{+}^{2}.$
All probabilities and expectations below are under these truncated measures.
For $v\in L\cup R$, define
$m_{v}:=\alpha+2\lambda\deg(v),\qquad m_{\min}:=\alpha+2\lambda\delta.$
Lemma 4.2 (Single-site conditional curvature).
For every $i$, $v$, and boundary condition on $L\cup R\setminus\{v\}$, the one-site conditional $\mu_{i}(f(v)\in\cdot\mid f_{-v})$ on an interval contained in [B,B][-B,B] is $m_{v}$-strongly log-concave.
Proof.
In the one-dimensional conditional potential, the base term contributes curvature $\alpha$. Each smoothing term $\lambda(X-c)^{2}$ from an incident edge contributes curvature $2\lambda$. Hinge-square terms are convex and add nonnegative curvature. The hard truncation only restricts the support interval and does not reduce strong convexity. ∎
Uniform control of boundary-margin failures.
A key step in adaptive spurious control is the following: when closing a fresh nonmatching edge with exactly one previously revealed endpoint, the support interval of $d=f(\ell)-f(r)$ must contain the target point $\gamma/2$. This can fail only if a previously revealed endpoint value lies extremely close to the hard truncation boundary $\pm B$. We show that such boundary-strip hits are uniformly unlikely under any transcript prefix, thanks to a deterministic inward drift created by the hinge terms when $\beta>\lambda$. Throughout this boundary-strip discussion, we impose the support condition $B\geq\gamma/2$.
Lemma 4.3 (Uniform inward drift near the bad boundary strips).
Fix $i\in[s]$, a vertex $v\in L\cup R$, and any boundary condition on $(L\cup R)\setminus\{v\}$. Let $X:=f(v)$ under the one-site conditional $\mu_{i}(\,\cdot\,\mid f_{-v})$, supported on an interval $J\subseteq[-B,B]$, with one-dimensional potential $h(\cdot)$ so that the conditional density is proportional to $\exp(-h(x))\mathbf{1}_{J}(x)$.
Assume $B\geq\gamma/2$ and $\beta>\lambda$. Then the following deterministic slope bounds hold (uniformly over the boundary condition):
  1. 1.
    If $v\in L$, then for every $x\in J\cap[B-\gamma/2,B]$,
    $h^{\prime}(x)\ \geq\ (\deg(v)-1)\cdot(\beta-\lambda)\gamma\ -\ (3\beta+\lambda)\gamma.$
  2. 2.
    If $v\in R$, then for every $x\in J\cap[-B,-B+\gamma/2]$,
    $h^{\prime}(x)\ \leq\ -(\deg(v)-1)\cdot(\beta-\lambda)\gamma\ +\ (3\beta+\lambda)\gamma.$
Proof.
Fix a vertex $v$ and condition on all other vertex-values (and on $I=i$), so that the one-site conditional law of $X_{v}$ has density proportional to $\exp(-h(x))$ on [B,B][-B,B] for some convex differentiable function $h$ (differentiable except at finitely-many hinge kinks, which is enough for the pointwise bounds below). Write $d:=\deg(v)$ and let $y_{u}:=f(u)\in[-B,B]$ denote the conditioned neighbor-values.
By definition, $h$ is the sum of: the vertex term $\frac{\alpha}{2}x^{2}$; and, for each incident edge, a smoothing term $\lambda(x-y_{u})^{2}$ plus a hinge-square term of coefficient $\beta$ whose form depends on whether the edge lies in $M_{i}$. Since $M_{i}$ is a matching, at most one incident edge of $v$ lies in $M_{i}$.
Case 1: $v\in L$ and $x\in[B-\gamma/2,B]$. For every nonmatching incident edge $(v,u)\notin M_{i}$ (with $u\in R$), the edge-contribution to $h$ equals
$\lambda(x-y_{u})^{2}+\beta(x-y_{u}+\gamma)_{+}^{2},$
hence its derivative is
$2\lambda(x-y_{u})+2\beta(x-y_{u}+\gamma)_{+}.$
On the strip $x\in[B-\gamma/2,B]$ we have $x-y_{u}+\gamma\geq(B-\gamma/2)-B+\gamma=\gamma/2$, so the hinge is active and $(x-y_{u}+\gamma)_{+}=x-y_{u}+\gamma\geq\gamma/2$. Also $x-y_{u}\geq(B-\gamma/2)-B=-\gamma/2$. Therefore each nonmatching incident edge contributes at least
$2\lambda(x-y_{u})+2\beta(x-y_{u}+\gamma)\ \geq\ 2\lambda(-\gamma/2)+2\beta(\gamma/2)\ =\ (\beta-\lambda)\gamma$
to $h^{\prime}(x)$.
If $v$ has a matching incident edge $(v,u_{0})\in M_{i}$, its edge-contribution is
$\lambda(x-y_{u_{0}})^{2}+\beta\big(\gamma-(x-y_{u_{0}})\big)_{+}^{2}=\lambda(x-y_{u_{0}})^{2}+\beta(y_{u_{0}}-x+\gamma)_{+}^{2},$
whose derivative is
$2\lambda(x-y_{u_{0}})-2\beta(y_{u_{0}}-x+\gamma)_{+}.$
On the same strip, $x-y_{u_{0}}\geq-\gamma/2$ so $2\lambda(x-y_{u_{0}})\geq-\lambda\gamma$, and $y_{u_{0}}-x+\gamma\leq B-(B-\gamma/2)+\gamma=3\gamma/2$ so $2\beta(y_{u_{0}}-x+\gamma)_{+}\geq-3\beta\gamma$. Thus the matching edge contributes at least $(3\beta+\lambda)\gamma$.
Finally, the vertex term contributes $\alpha x\geq 0$ on this upper strip since $x\geq B-\gamma/2\geq 0$ under $B\geq\gamma/2$. Summing the contributions, we obtain
$h^{\prime}(x)\ \geq\ (\deg(v)-1)\cdot(\beta-\lambda)\gamma\ -\ (3\beta+\lambda)\gamma,$
which is the first claimed bound.
Case 2: $v\in R$ and $x\in[-B,-B+\gamma/2]$. Now each nonmatching incident edge has the form $(u,v)\notin M_{i}$ with $u\in L$, and contributes
$\lambda(y_{u}-x)^{2}+\beta(y_{u}-x+\gamma)_{+}^{2}$
to $h$. Its derivative is
$2\lambda(x-y_{u})-2\beta(y_{u}-x+\gamma)_{+}.$
On the strip $x\in[-B,-B+\gamma/2]$ we have $y_{u}-x+\gamma\geq(-B)-(-B+\gamma/2)+\gamma=\gamma/2$, so the hinge is active and $(y_{u}-x+\gamma)_{+}\geq\gamma/2$. Also $x-y_{u}\leq(-B+\gamma/2)-(-B)=\gamma/2$, so $2\lambda(x-y_{u})\leq\lambda\gamma$. Hence each nonmatching incident edge contributes at most
$2\lambda(x-y_{u})-2\beta(y_{u}-x+\gamma)\ \leq\ \lambda\gamma-\beta\gamma\ =\ -(\beta-\lambda)\gamma$
to $h^{\prime}(x)$.
If $v$ has a matching incident edge $(u_{0},v)\in M_{i}$, its edge-contribution is
$\lambda(y_{u_{0}}-x)^{2}+\beta\big(\gamma-(y_{u_{0}}-x)\big)_{+}^{2}=\lambda(y_{u_{0}}-x)^{2}+\beta(\gamma-y_{u_{0}}+x)_{+}^{2},$
whose derivative is
$2\lambda(x-y_{u_{0}})+2\beta(\gamma-y_{u_{0}}+x)_{+}.$
On our strip, $x-y_{u_{0}}\leq\gamma/2$ so $2\lambda(x-y_{u_{0}})\leq\lambda\gamma$, and $\gamma-y_{u_{0}}+x\leq\gamma-(-B)+(-B+\gamma/2)=3\gamma/2$ so the hinge term is at most $3\beta\gamma$. Thus the matching edge contributes at most $(3\beta+\lambda)\gamma$.
Finally, the vertex term contributes $\alpha x\leq 0$ on this lower strip since $x\leq-B+\gamma/2\leq 0$ under $B\geq\gamma/2$, so it only helps the upper bound. Summing, we obtain
$h^{\prime}(x)\ \leq\ -(\deg(v)-1)\cdot(\beta-\lambda)\gamma\ +\ (3\beta+\lambda)\gamma,$
which is the second claimed bound. ∎
Lemma 4.4 (Uniform boundary-strip probability bound).
Assume $B\geq\gamma/2$, $\beta\geq 2\lambda$, $\delta\geq 10$, and moreover $\alpha\leq(\lambda+\beta)\delta$. Then there exist absolute constants $c_{\mathrm{bdry}},C_{\mathrm{bdry}}>0$ such that for every $i\in[s]$, every vertex $v$ with $\deg(v)\geq\delta$, and every boundary condition on $(L\cup R)\setminus\{v\}$,
$\Pr_{\mu_{i}(\cdot\mid f_{-v})}\bigl[X\in[B-\gamma/2,B]\bigr]\ \leq\ C_{\mathrm{bdry}}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}\cdot\exp\!\left(-c_{\mathrm{bdry}}\cdot\delta\cdot\frac{(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2}\right)\qquad(v\in L),$
and
$\Pr_{\mu_{i}(\cdot\mid f_{-v})}\bigl[X\in[-B,-B+\gamma/2]\bigr]\ \leq\ C_{\mathrm{bdry}}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}\cdot\exp\!\left(-c_{\mathrm{bdry}}\cdot\delta\cdot\frac{(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2}\right)\qquad(v\in R).$
Proof.
We prove the first claim (for $v\in L$); the second is analogous by left/right symmetry.
Fix $v\in L$ and fix any conditioning on $\{f(u):u\neq v\}$ and on $I=i$, so that $X:=f(v)$ has density proportional to $\exp(-h(x))$ on [B,B][-B,B], for a convex function $h$. By Section 4.1, $h$ is $m_{v}$-strongly convex, i.e. $h^{\prime\prime}\geq m_{v}$ (in the sense of distributions). Moreover, since each quadratic term contributes curvature $2\lambda$ and each hinge-square term contributes curvature at most $2\beta$ when active, we also have a uniform upper curvature bound
$h^{\prime\prime}\ \leq\ M_{v}\ :=\ \alpha+2(\lambda+\beta)\deg(v)\qquad\text{on }(-B,B).$ (6)
Let $x^{\star}$ be a minimizer of $h$ on [B,B][-B,B] (so $h^{\prime}(x^{\star})=0$ if $x^{\star}\in(-B,B)$). Let $a:=B-\gamma/2$.
Step 1: a uniform slope at the strip. By Section 4.1, for all $x\in[a,B]$,
$h^{\prime}(x)\ \geq\ (\deg(v)-1)(\beta-\lambda)\gamma-(3\beta+\lambda)\gamma\ =:\ A.$
Under $\deg(v)\geq\delta\geq 10$ and $\beta\geq 2\lambda$, this satisfies $A>0$ and in fact $A\geq c_{1}\,\deg(v)(\beta-\lambda)\gamma$ for a universal constant $c_{1}>0$.
Step 2: mode separation. Step 1 shows that $h^{\prime}(x)\geq A>0$ for all $x\in[a,B]$. Therefore any minimizer $x^{\star}$ of $h$ on [B,B][-B,B] lies in [B,a][-B,a] (in particular $x^{\star}\neq B$). If $x^{\star}\in(-B,a]$, then $h^{\prime}(x^{\star})=0$, and since $h^{\prime}$ is $M_{v}$-Lipschitz by (6),
$A\ \leq\ h^{\prime}(a)-h^{\prime}(x^{\star})\ \leq\ M_{v}\,(a-x^{\star}),\qquad\text{hence}\qquad a-x^{\star}\ \geq\ A/M_{v}.$
If instead $x^{\star}=-B$, then $a-x^{\star}=a+B=2B-\gamma/2$. In this case the same bound still holds because
$\frac{A}{M_{v}}\ \leq\ \frac{\deg(v)(\beta-\lambda)\gamma}{2(\lambda+\beta)\deg(v)}\ \leq\ \frac{\gamma}{2}\ \leq\ a+B$
using $A\leq\deg(v)(\beta-\lambda)\gamma$, $M_{v}\geq 2(\lambda+\beta)\deg(v)$, and $B\geq\gamma/2$. Thus in all cases,
$a-x^{\star}\ \geq\ A/M_{v}.$ (7)
Step 3: a potential gap at the strip. Strong convexity ($h^{\prime\prime}\geq m_{v}$) implies $h(x)\geq h(x^{\star})+\tfrac{m_{v}}{2}(x-x^{\star})^{2}$ for all $x$. In particular, by (7),
$h(a)-h(x^{\star})\ \geq\ \frac{m_{v}}{2}(a-x^{\star})^{2}\ \geq\ \frac{m_{v}A^{2}}{2M_{v}^{2}}.$ (8)
Step 4: compare numerator and denominator. For $x\in[a,B]$, convexity gives $h(x)\geq h(a)+h^{\prime}(a)(x-a)\geq h(a)+A(x-a)$, so
$\int_{a}^{B}e^{-h(x)}\,dx\ \leq\ e^{-h(a)}\int_{0}^{\infty}e^{-At}\,dt\ =\ \frac{e^{-h(a)}}{A}.$ (9)
For the denominator $Z:=\int_{-B}^{B}e^{-h(x)}dx$, we lower bound the normalizing constant by integrating over a small neighborhood of the minimizer. Since $X\in[-B,B]$ and all conditioned neighbor values lie in [B,B][-B,B], every term in $h^{\prime}(x)$ is uniformly bounded on [B,B][-B,B]: the vertex term contributes at most $\alpha B$ in magnitude, each quadratic edge term contributes at most $4\lambda B$, and each hinge-squared term contributes at most $2\beta(2B+\gamma)$. Therefore
$|h^{\prime}(x)|\ \leq\ L_{v}\ :=\ \alpha B+\deg(v)\bigl(4\lambda B+2\beta(2B+\gamma)\bigr)\qquad\text{for all }x\in(-B,B),$ (10)
and hence $h$ is $L_{v}$-Lipschitz on [B,B][-B,B]. Let $r:=1/(2L_{v})$ and $I:=([x^{\star}-r,\,x^{\star}+r]\cap[-B,B])$. Then $|I|\geq r$ and $h(x)\leq h(x^{\star})+1/2$ for all $x\in I$, so
[…]
Combining (9)–(11) and using (8) gives
$\Pr[X\in[a,B]\mid\text{conditioning}]\ \leq\ \frac{2L_{v}}{A}\,e^{1/2}\,\exp\!\Bigl(-\frac{m_{v}A^{2}}{2M_{v}^{2}}\Bigr).$
Step 5: bound the prefactor explicitly. Using $m_{v}=\alpha+2\lambda\deg(v)\geq 2\lambda\deg(v)$ and, since $\deg(v)\geq\delta$ and we assume $\alpha\leq(\lambda+\beta)\delta$, we have $M_{v}\leq\alpha+2(\lambda+\beta)\deg(v)\leq 3(\lambda+\beta)\deg(v)$, together with $A\geq c_{1}\deg(v)(\beta-\lambda)\gamma$, we obtain
$\frac{m_{v}A^{2}}{2M_{v}^{2}}\ \geq\ c_{2}\cdot\deg(v)\cdot\frac{(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2}\ \geq\ c_{2}\cdot\delta\cdot\frac{(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2},$
for a universal constant $c_{2}>0$.
For the prefactor, recall $L_{v}=\alpha B+\deg(v)\bigl(4\lambda B+2\beta(2B+\gamma)\bigr)$ from (10) and $A\geq c_{1}\deg(v)(\beta-\lambda)\gamma$. Since $\deg(v)\geq 1$ and $\beta\geq 2\lambda$, we have
$\frac{2L_{v}}{A}\ \leq\ C_{3}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}$
for a universal constant $C_{3}>0$. Combining this with the bound at the end of Step 4 yields the lemma with $C_{\mathrm{bdry}}:=e^{1/2}C_{3}$ and (after possibly decreasing $c_{2}$ by an absolute factor) $c_{\mathrm{bdry}}:=c_{2}$. ∎
Lemma 4.5 (Score expectation equals boundary leakage (1D integration-by-parts)).
Let $X$ have density $p(x)=Z^{-1}e^{-h(x)}\mathbf{1}_{[a,b]}(x)$ where $h$ is absolutely continuous on $(a,b)$. Then
$\mathbb{E}[h^{\prime}(X)]\ =\ \frac{e^{-h(b)}-e^{-h(a)}}{\int_{a}^{b}e^{-h(x)}\,dx}.$
Proof.
Differentiate $e^{-h(x)}$ and integrate over $[a,b]$:
$\int_{a}^{b}h^{\prime}(x)e^{-h(x)}\,dx\;=\;-\int_{a}^{b}\frac{d}{dx}\bigl(e^{-h(x)}\bigr)\,dx\;=\;e^{-h(b)}-e^{-h(a)}.$
Dividing by $Z=\int_{a}^{b}e^{-h(x)}\,dx$ gives the claim. ∎
4.2 Edge-tail bounds and admissible parameter set
This subsection derives edge-tail bounds under the truncated model and states the parameter conditions needed for these bounds.
Probabilistic control of the external drift.
Fix an edge $e=(\ell,r)$ and the difference coordinate $d=f(\ell)-f(r)$. Informally, all constraints incident to $\ell$ or $r$ other than the designated edge $e$ create an external “force” on $d$; we call this contribution the (external) drift. In formulas, this drift is the term $R_{e,t}(d)$, and $R^{\prime}_{e,t}(d):=\frac{d}{dd}R_{e,t}(d)$ is its directional effect on the $d$-coordinate. A precise definition is given by the decomposition (21) in the proof of Section 4.2. In our analysis, we prove a high-probability bound at the two points $d_{0}\in\{\pm\gamma/2\}$ that appear in the edge-tail argument. Fix a drift tolerance $u>0$ and define
$p_{\mathrm{drift}}(u)\;:=\;2\exp\!\left(-\,c_{\mathrm{drift}}\cdot\frac{m_{\min}\,u^{2}}{(\lambda+\beta)^{2}\,\Delta}\right),$ (12)
where $c_{\mathrm{drift}}>0$ is an absolute constant.
Additional graph-theoretic assumption.
The next lemma (and only this lemma) uses an additional graph-theoretic property of the base graph: we assume that the underlying undirected graph of $U$ is $C_{4}$-free (has no $4$-cycles). This hypothesis is satisfied by the explicit family constructed in Section 4.7 below. All subsequent uses of $p_{\mathrm{drift}}(u)$ and the edge-tail bounds depend on this hypothesis.
Lemma 4.6 (External drift tail bound for $C_{4}$-free graphs).
Assume that the underlying undirected graph of $U$ contains no $4$-cycles (equivalently: $U$ is $C_{4}$-free). Fix any time $t$ and any edge $e=(\ell,r)$. Then for each $d_{0}\in\{\pm\gamma/2\}$ and every conditioning on $(I=i,T_{<t})$,
$\Pr\bigl[\,|R^{\prime}_{e,t}(d_{0})-\mathbb{E}[R^{\prime}_{e,t}(d_{0})\mid\mathcal{F},f(\ell),f(r)]|>u\,\bigm|\,I=i,\;T_{<t}\bigr]\ \leq\ p_{\mathrm{drift}}(u).$
Proof.
Fix $t,e=(\ell,r)$ and $d_{0}\in\{\pm\gamma/2\}$. Let $\mathcal{F}$ denote the $\sigma$-field generated by $(I=i,T_{<t})$ together with the values of all vertices outside the one-neighborhood of $\{\ell,r\}$. Condition on $\mathcal{F}$ throughout this proof.
Write
$\Gamma:=\bigl(N(\ell)\cup N(r)\bigr)\setminus\{\ell,r\}.$
[…]
Next, evaluate the external drift $R^{\prime}_{e,t}(d)$ at $d=d_{0}$. By definition, $R_{e,t}(d)$ collects the contributions to the one-dimensional potential on the difference coordinate $d$ from all edges incident to $\ell$ or $r$ except the designated edge $e$ itself. Each such incident edge involves exactly one neighbor $v\in\Gamma$ together with one of the endpoints. For the fixed discrete data $(i,t,e,d_{0})$ (and the edge-type of the incident constraint), the contribution of that incident edge to $R^{\prime}_{e,t}(d_{0})$ can be written as a function of the neighbor value alone. Hence there exist measurable functions $G_{v}:\mathbb{R}\to\mathbb{R}$ such that
$R^{\prime}_{e,t}(d_{0})\;=\;\sum_{v\in\Gamma}G_{v}\!\bigl(f(v)\bigr).$
We stress that, once the discrete data $(i,t,e,d_{0})$ are fixed and we condition on the endpoint values $(f(\ell),f(r))$, each incident-edge contribution becomes a deterministic function of the single neighbor spin $f(v)$. (In the open-edge case this one-variable function may depend parametrically on the already revealed endpoint value, since the probe location of the moving endpoint at $d=d_{0}$ depends on it, but this dependence is frozen under the conditioning on $T_{<t}$.) Conditioning on $(f(\ell),f(r))$ is therefore used only to guarantee that $\{f(v)\}_{v\in\Gamma}$ are independent. Moreover, each $G_{v}$ is $(\lambda+\beta)$-Lipschitz: for quadratic terms $\lambda(f(u)-f(v))^{2}$ the derivative contribution is linear with slope $\lambda$, and for hinge-square terms $\beta(\,\cdot\,)_{+}^{2}$ the derivative contribution has slope at most $\beta$. Thus, for every $v\in\Gamma$,
$\mathrm{Lip}(G_{v})\ \leq\ \lambda+\beta.$ (13)
Now fix $v\in\Gamma$ and consider the (conditional) one-dimensional law of $X_{v}:=f(v)$. By Section 4.1, this one-site conditional distribution is $m_{v}$-strongly log-concave on [B,B][-B,B] with $m_{v}\geq m_{\min}$. A standard consequence (e.g. 1D Bakry–Émery/Herbst) is that for any $L$-Lipschitz function $g$,
$\mathbb{E}\bigl[e^{\theta(g(X_{v})-\mathbb{E}g(X_{v}))}\bigr]\ \leq\ \exp\!\left(\frac{\theta^{2}L^{2}}{2m_{\min}}\right)\qquad\text{for all }\theta\in\mathbb{R}.$ (14)
Applying (14) to $g=G_{v}$ with $L=\lambda+\beta$ gives that $G_{v}(X_{v})-\mathbb{E}G_{v}(X_{v})$ is subgaussian with proxy $(\lambda+\beta)^{2}/m_{\min}$.
Since $\{X_{v}\}_{v\in\Gamma}$ are independent conditional on $\mathcal{F}$ and $(f(\ell),f(r))$, we can sum the mgf bounds and obtain
$\mathbb{E}\!\left[\exp\!\Bigl(\theta\bigl(R^{\prime}_{e,t}(d_{0})-\mathbb{E}[R^{\prime}_{e,t}(d_{0})]\bigr)\Bigr)\,\Bigm|\,\mathcal{F},f(\ell),f(r)\right]\ \leq\ \exp\!\left(\frac{\theta^{2}(\lambda+\beta)^{2}|\Gamma|}{2m_{\min}}\right).$
Finally, $|\Gamma|\leq 2\Delta$ and $\frac{1}{m_{\min}}\leq\frac{\Delta}{m_{\min}}\cdot\frac{1}{\Delta}$, so by Chernoff,
$\Pr\bigl[\,|R^{\prime}_{e,t}(d_{0})-\mathbb{E}R^{\prime}_{e,t}(d_{0})|>u\,\bigm|\,\mathcal{F},f(\ell),f(r)\bigr]\ \leq\ 2\exp\!\left(-\,c\cdot\frac{m_{\min}\,u^{2}}{(\lambda+\beta)^{2}\,\Delta}\right).$
Absorbing constants and removing the extra conditioning (tower property) yields (12). ∎
Remark (centering vs. absolute drift). Section 4.2 gives concentration only around the conditional mean. To derive the one-sided bounds on $R^{\prime}_{e,t}(\pm\gamma/2)$ needed in the edge-tail argument, we combine Section 4.2 with the mean-bad events (17); see Section 4.2.
Admissible parameter conditions.
We impose the following support condition (strong enough to fix the sign on the boundary strips used below):
[…]
Given the drift tolerance $u$, we use the following separation condition:
$\beta\gamma\ \geq\ \alpha\!\left(B+\frac{\gamma}{2}\right)+\lambda\gamma+u.$ (16)
Mean and drift-bad events.
Section 4.2 controls fluctuations about the conditional mean. For the edge-tail argument we only need one-sided control on $R^{\prime}_{e,t}(d_{0})$: an upper tail at $d_{0}=+\gamma/2$ (used for matching edges) and a lower tail at $d_{0}=-\gamma/2$ (used for nonmatching edges). We therefore isolate the conditional mean as a separate failure event.
For $u>0$ define
$\mathsf{MeanBad}^{+}_{e,t}$ $=\left\{\mathbb{E}\bigl[R^{\prime}_{e,t}(\gamma/2)\mid\mathcal{F},f(\ell),f(r)\bigr]>\frac{u}{2}\right\},$ (17)
$\mathsf{MeanBad}^{-}_{e,t}$ $=\left\{\mathbb{E}\bigl[R^{\prime}_{e,t}(-\gamma/2)\mid\mathcal{F},f(\ell),f(r)\bigr]<-\frac{u}{2}\right\}.$
For an edge $e=(\ell,r)$ and round $t$, let $\mathsf{Open}_{e,t}$ be the event that the transcript prefix $T_{<t}$ has revealed at most one of the endpoint values $f(\ell),f(r)$ (equivalently, the edge is not yet fully revealed before round $t$). In all applications of the drift bounds below, we only invoke them for such “open” edges (indeed, Section 4.2(1) corresponds to $t=1$ so $T_{<t}$ is empty, while in Section 4.2(2) exactly one endpoint is revealed by assumption). We therefore define the mean-bad probability only over this regime:
$p_{\mathrm{mean}}(u):=\sup_{i,t,e}\Pr\bigl(\mathsf{MeanBad}^{+}_{e,t}\;\text{or}\;\mathsf{MeanBad}^{-}_{e,t}\mid I=i,T_{<t},\ \mathsf{Open}_{e,t}\bigr).$ (18)
Lemma 4.7 (From drift-tail to one-sided drift bounds).
Fix $u>0$. For any $i,t,e=(\ell,r)$,
$\Pr\bigl(R^{\prime}_{e,t}(\gamma/2)>u\mid I=i,T_{<t},\ \mathsf{Open}_{e,t}\bigr)\;\leq\;p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u),$
and
$\Pr\bigl(R^{\prime}_{e,t}(-\gamma/2)<-u\mid I=i,T_{<t},\ \mathsf{Open}_{e,t}\bigr)\;\leq\;p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u).$
Proof.
We prove the first inequality; the second is identical with signs reversed. Condition on $\mathcal{F},f(\ell),f(r)$. On the complement of $\mathsf{MeanBad}_{e,t}^{+}$ we have $\mathbb{E}[R^{\prime}_{e,t}(\gamma/2)\mid\mathcal{F},f(\ell),f(r)]\leq u/2$. Hence the event $\{R^{\prime}_{e,t}(\gamma/2)>u\}$ implies
$R^{\prime}_{e,t}(\gamma/2)-\mathbb{E}[R^{\prime}_{e,t}(\gamma/2)\mid\mathcal{F},f(\ell),f(r)]>u/2.$
Define
$\Xi_{e,t}:=R^{\prime}_{e,t}(\gamma/2)-\mathbb{E}[R^{\prime}_{e,t}(\gamma/2)\mid\mathcal{F},f(\ell),f(r)],\qquad\mathcal{C}:=\{I=i,\ T_{<t},\ \mathsf{Open}_{e,t}\}.$
Taking probabilities and union bounding gives
$\Pr\bigl(R^{\prime}_{e,t}(\gamma/2)>u\mid\mathcal{C}\bigr)$ $\leq\Pr\bigl(\mathsf{MeanBad}_{e,t}^{+}\mid\mathcal{C}\bigr)$
$\quad+\Pr\bigl(|\Xi_{e,t}|>u/2\mid\mathcal{C}\bigr).$
The first term is bounded by $p_{\mathrm{mean}}(u)$ by definition (18). The second term is bounded by $p_{\mathrm{drift}}(u/2)$ by definition (12) and Section 4.2. ∎
We now return to the edge-tail argument. For the one-dimensional edge calculations below, define
$m_{\mathrm{edge}}\;:=\;\frac{\alpha}{2}+2\lambda,$ (19)
the uniform lower bound on the curvature coming from the quadratic (pinning and smoothing) part on a single directed edge.
Define
$p_{\mathrm{match}}$ $=\exp(-m_{\mathrm{edge}}\gamma^{2}/8)+p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u),$ (20)
$p_{\mathrm{spur}}$ $=\exp(-m_{\mathrm{edge}}\gamma^{2}/8)+p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u).$
Remark. We keep separate notation for $p_{\mathrm{match}}$ and $p_{\mathrm{spur}}$ (despite the identical expression in (20)) to emphasize that they control different events: non-violation on a matching edge vs. spurious violation on a nonmatching edge under adaptive edge-closing.
Lemma 4.8 (Edge-tail bounds).
Assume (15) and (16). Fix any $i\in[s]$.
  1. 1.
    If $e=(\ell,r)\in M_{i}$, then
    $\Pr_{f\sim\mu_{i}}\bigl[f(\ell)\leq f(r)\bigr]\ \leq\ p_{\mathrm{match}}.$
  2. 2.
    If $e=(\ell,r)\notin M_{i}$ and at time $t$ exactly one endpoint value is revealed so that $\gamma/2\in I_{e,t}$, then
    $\Pr_{f\sim\mu_{i}}\bigl[f(\ell)>f(r)\,\bigm|\,T_{<t}\bigr]\ \leq\ p_{\mathrm{spur}}.$
Proof.
We prove both parts by reducing to a one-dimensional log-concave distribution on the difference coordinate and then applying strong convexity plus a sign-of-derivative argument. We condition on the drift being well-controlled at the relevant point $d_{0}\in\{\pm\gamma/2\}$. Define
$\mathsf{DriftGood}^{+}_{e,t}:=\Bigl\{\bigl|R^{\prime}_{e,t}(\gamma/2)-\mathbb{E}[R^{\prime}_{e,t}(\gamma/2)\mid\mathcal{F},f(\ell),f(r)]\bigr|\leq u/2\Bigr\}\cap\bigl(\mathsf{MeanBad}^{+}_{e,t}\bigr)^{c},$
and
$\mathsf{DriftGood}^{-}_{e,t}:=\Bigl\{\bigl|R^{\prime}_{e,t}(-\gamma/2)-\mathbb{E}[R^{\prime}_{e,t}(-\gamma/2)\mid\mathcal{F},f(\ell),f(r)]\bigr|\leq u/2\Bigr\}\cap\bigl(\mathsf{MeanBad}^{-}_{e,t}\bigr)^{c}.$
On $\mathsf{DriftGood}^{+}_{e,t}$ we have $R^{\prime}_{e,t}(\gamma/2)\leq u$, and on $\mathsf{DriftGood}^{-}_{e,t}$ we have $R^{\prime}_{e,t}(-\gamma/2)\geq-u$. Moreover, by Section 4.2,
$\Pr\bigl((\mathsf{DriftGood}^{+}_{e,t})^{c}\mid I=i,T_{<t},\ \mathsf{Open}_{e,t}\bigr)$ $\leq p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u),$
$\Pr\bigl((\mathsf{DriftGood}^{-}_{e,t})^{c}\mid I=i,T_{<t},\ \mathsf{Open}_{e,t}\bigr)$ $\leq p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u).$
In part (1) we may take $t=1$ so the prefix transcript $T_{<t}$ is empty, and hence $\mathsf{Open}_{e,t}$ holds automatically. In part (2), $\mathsf{Open}_{e,t}$ holds by the assumption that exactly one endpoint is revealed.
Part (1): $e\in M_{i}$. Let $d:=f(\ell)-f(r)$. Condition on all vertex values except $(f(\ell),f(r))$ and change variables to $(z,d)$ via $f(\ell)=z+\tfrac{d}{2}$ and $f(r)=z-\tfrac{d}{2}$. Integrating out $z$ yields a one-dimensional density on $d\in[-2B,2B]$ of the form
$\pi(d)\ \propto\ \exp\!\bigl(-W_{e,t}(d)\bigr),$
where $W_{e,t}$ is $m_{\mathrm{edge}}$-strongly convex and can be decomposed as
$W_{e,t}(d)\;=\;U_{e,t}(d)\;+\;\beta\,\Psi_{e,i}(d)\;+\;R_{e,t}(d).$ (21)
Here $U_{e,t}$ is the quadratic part (pinning and smoothing) involving the endpoints and the edge $e$, $\Psi_{e,i}(d)=(\gamma-d)_{+}^{2}$ is the designated hinge-square term for a matching edge, and $R_{e,t}$ collects all other incident terms.
Let $d_{\star}$ be a minimizer of $W_{e,t}$. If $W_{e,t}^{\prime}(\gamma/2)\leq 0$, then by convexity $d_{\star}\geq\gamma/2$. On the event $\mathsf{DriftGood}^{+}_{e,t}$ we have
$W^{\prime}_{e,t}(\gamma/2)\;\leq\;\underbrace{U^{\prime}_{e,t}(\gamma/2)}_{\leq\ \alpha(B+\gamma/2)+\lambda\gamma}\;+\;\underbrace{\beta\,\Psi^{\prime}_{e,i}(\gamma/2)}_{=-\beta\gamma}\;+\;\underbrace{R^{\prime}_{e,t}(\gamma/2)}_{\leq\ u}\ \leq\ 0,$
where we used (16) in the last step. Thus $d_{\star}\geq\gamma/2$ on $\mathsf{DriftGood}^{+}_{e,t}$. Strong convexity then yields the left-tail bound
$\Pr[d\leq 0\mid\mathsf{DriftGood}^{+}_{e,t}]\ \leq\ \exp\!\left(-\frac{m_{\mathrm{edge}}\gamma^{2}}{8}\right).$
Finally, since $(\mathsf{DriftGood}^{+}_{e,t})^{c}=\mathsf{MeanBad}^{+}_{e,t}\cup\mathsf{DriftBad}^{+}_{e,t}(u/2)$, a union bound gives
$\Pr[d\leq 0]\ \leq\ \exp\!\left(-\frac{m_{\mathrm{edge}}\gamma^{2}}{8}\right)+p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u)=p_{\mathrm{match}}.$
Part (2): $e\notin M_{i}$ and $\gamma/2\in I_{e,t}$. At time $t$ exactly one endpoint is revealed, say $f(\ell)=a\in[-B,B]$. Then d=af(r)d=a-f(r) ranges over the interval $I_{e,t}=[a-B,a+B]$, which contains $\gamma/2$ by assumption. As in part (1), conditioning on all other vertices yields a one-dimensional density $\pi(d)\propto\exp(-W_{e,t}(d))$ supported on $I_{e,t}$, with $W_{e,t}$ again $m_{\mathrm{edge}}$-strongly convex. Now $\Psi_{e,i}(d)=(d+\gamma)_{+}^{2}$ since $e\notin M_{i}$. Let $d_{\star}$ be a minimizer of $W_{e,t}$. If $W^{\prime}_{e,t}(-\gamma/2)\geq 0$, then by convexity $d_{\star}\leq-\gamma/2$. On the event $\mathsf{DriftGood}^{-}_{e,t}$ we have
$W^{\prime}_{e,t}(-\gamma/2)\;\geq\;\underbrace{U^{\prime}_{e,t}(-\gamma/2)}_{\geq\ -\alpha(B+\gamma/2)-\lambda\gamma}\;+\;\underbrace{\beta\,\Psi^{\prime}_{e,i}(-\gamma/2)}_{=+\beta\gamma}\;+\;\underbrace{R^{\prime}_{e,t}(-\gamma/2)}_{\geq\ -u}\ \geq\ 0,$
again by (16). Hence $d_{\star}\leq-\gamma/2$ on $\mathsf{DriftGood}^{-}_{e,t}$, and strong convexity gives
$\Pr[d>0\mid\mathsf{DriftGood}^{-}_{e,t}]\ \leq\ \exp\!\left(-\frac{m_{\mathrm{edge}}\gamma^{2}}{8}\right).$
Finally, since $(\mathsf{DriftGood}^{-}_{e,t})^{c}=\mathsf{MeanBad}^{-}_{e,t}\cup\mathsf{DriftBad}^{-}_{e,t}(u/2)$, a union bound gives
$\Pr[d>0\mid T_{<t}]$ $\leq\Pr[d>0\mid\mathsf{DriftGood}^{-}_{e,t}]+\Pr[(\mathsf{DriftGood}^{-}_{e,t})^{c}\mid T_{<t}]$
$\leq\exp\!\left(-\frac{m_{\mathrm{edge}}\gamma^{2}}{8}\right)+p_{\mathrm{drift}}(u/2)+p_{\mathrm{mean}}(u)$
$=p_{\mathrm{spur}}.$
as claimed. ∎
4.3 Farness and control of spurious queried violations
This subsection upgrades the edge-tail estimates to global farness and bounds the probability of adaptive spurious witnesses.
Define
$\mathrm{GoodMatch}_{i}:=\Big\{\#\{e\in M_{i}:e\text{ not violated}\}\leq|M_{i}|/100\Big\}.$
Lemma 4.9 (Matching-good event has high probability).
If $p_{\mathrm{match}}\leq 10^{-4}$, then for every $i$,
$\Pr_{f\sim\mu_{i}}[\mathrm{GoodMatch}_{i}]\geq 0.99.$
Proof.
Let $B_{i}$ be the number of non-violated edges in $M_{i}$. By Section 4.2(1),
$\mathbb{E}[B_{i}]\leq|M_{i}|p_{\mathrm{match}}\leq|M_{i}|/10^{4}.$
Markov at threshold $|M_{i}|/100$ gives
$\Pr[B_{i}>|M_{i}|/100]\leq 0.01.$
Lemma 4.10 (Farness with high probability).
Under Section 4.3’s assumption, for each $i$,
$\Pr_{f\sim\mu_{i}}\big[d_{\mathrm{mon}}(f)\geq 2(\varepsilon_{0}/200)n\big]\geq 0.99.$
Proof.
On $\mathrm{GoodMatch}_{i}$, at least $99/100$ of $M_{i}$ is violated, so violated edges contain a matching of size
$(99/100)|M_{i}|\geq(99/100)\varepsilon_{0}n_{0}=\frac{99}{200}\varepsilon_{0}n\geq 2(\varepsilon_{0}/200)n.$
Apply Section 3.3.1. ∎
Lemma 4.11 (Adaptive spurious-witness bound).
Fix a deterministic adaptive one-sided tester making $q\leq Q$ queries. For each round $t$, let $\mathrm{NewEdges}_{t}(T)$ be the set of newly closed edges in $E\setminus M_{I}$ (i.e., edges in $E\setminus M_{I}$ whose second endpoint is queried at round $t$). For $e\in\mathrm{NewEdges}_{t}(T)$, let $\mathcal{I}_{e,t}$ be the support interval of $d=f(\ell)-f(r)$ under $(I,T_{<t})$. Define
$\mathsf{SpurQry}:=\{\exists\text{ queried violated edge in }E\setminus M_{I}\}.$
Also define the boundary-margin failure event
$\mathsf{BdryQry}:=\{\exists t,\exists e\in\mathrm{NewEdges}_{t}(T)\text{ with }-\gamma/2\notin\mathcal{I}_{e,t}\}.$
If
$p_{\mathrm{spur}}\leq\frac{1}{100Q^{2}},$
then
$\Pr[\mathsf{SpurQry}]\leq\Pr[\mathsf{BdryQry}]+0.01.$
Proof.
Condition on $I=i$. Let $\mathcal{E}_{\mathrm{qry}}(T)$ be edges whose two endpoints are queried. Since at most $q$ vertices are queried,
$|\mathcal{E}_{\mathrm{qry}}(T)|\leq q^{2}/4.$
Expose queried vertices sequentially. For each round $t$, let
$G_{t}:=\{e\in\mathrm{NewEdges}_{t}(T):-\gamma/2\in\mathcal{I}_{e,t}\},$
and $B_{t}:=\mathrm{NewEdges}_{t}(T)\setminus G_{t}$. By Section 4.2 (Case B, nonmatching part),
$\Pr[\exists\text{ violated edge in }G_{t}\mid T_{<t},I=i]\leq|G_{t}|\,p_{\mathrm{spur}}.$
Hence
$\Pr[\exists\text{ violated edge in }\mathrm{NewEdges}_{t}(T)\mid T_{<t},I=i]\leq\mathbf{1}_{\{B_{t}\neq\emptyset\}}+|G_{t}|\,p_{\mathrm{spur}}.$
Summing over $t$, taking expectations, and using $\sum_{t}|G_{t}|\leq|\mathcal{E}_{\mathrm{qry}}(T)|\leq q^{2}/4$:
$\Pr[\mathsf{SpurQry}\mid I=i]$ $\leq\Pr[\mathsf{BdryQry}\mid I=i]+\frac{q^{2}}{4}p_{\mathrm{spur}}$
$\leq\Pr[\mathsf{BdryQry}\mid I=i]+\frac{Q^{2}}{4}\cdot\frac{1}{100Q^{2}}$
$\leq 0.01+\Pr[\mathsf{BdryQry}\mid I=i].$
Average over $i$. ∎
Lemma 4.12 (Strategy-independent bound on the boundary-margin failure).
Assume (15), $\beta\geq 2\lambda$, $\delta\geq 10$, and moreover $\alpha\leq(\lambda+\beta)\delta$. Let $\mathcal{A}$ be any deterministic adaptive strategy making $q\leq Q$ queries, and recall $\eta_{\partial,\mathcal{A}}:=\Pr_{\mathcal{A}}[\mathsf{BdryQry}]$. Then
$\eta_{\partial,\mathcal{A}}\ \leq\ Q\cdot C_{\mathrm{bdry}}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}\cdot\exp\!\left(-c_{\mathrm{bdry}}\cdot\delta\cdot\frac{(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2}\right),$
where $c_{\mathrm{bdry}},C_{\mathrm{bdry}}>0$ are the absolute constants from Section 4.1. In particular, choosing $\gamma=C\sqrt{\log(100Q^{2})}$ with a sufficiently large absolute constant $C$ makes $\eta_{\partial,\mathcal{A}}\leq 0.01$ uniformly over all deterministic strategies.
Proof.
We first relate $\mathsf{BdryQry}$ to hitting a narrow boundary strip. If $\mathsf{BdryQry}$ occurs, then for some round $t$ there exists a newly closed nonmatching edge $e=(\ell,r)\in\mathrm{NewEdges}_{t}(T)$ such that $\gamma/2\notin\mathcal{I}_{e,t}$. In Case B of Section 4.2, exactly one endpoint was previously revealed, with value $a$.
If the revealed endpoint is $\ell\in L$, then the support interval is $\mathcal{I}_{e,t}=[a-B,a+B]$, so $\gamma/2\notin\mathcal{I}_{e,t}$ implies $a>B-\gamma/2$.
If the revealed endpoint is $r\in R$, then $\mathcal{I}_{e,t}=[-B-a,B-a]$, so $\gamma/2\notin\mathcal{I}_{e,t}$ implies $a<-B+\gamma/2$ (the alternative inequality $\gamma/2>B-a$ would force $a>B+\gamma/2$, impossible since $a\leq B$).
Thus $\mathsf{BdryQry}$ implies that at some query round, the revealed value fell in the corresponding boundary strip $B-\gamma/2,B]$ (for $L$-queries) or $B,-B+\gamma/2]$ (for $R$-queries).
Now fix any round $t$ and condition on the prefix $T_{<t}$. The answer at round $t$ is drawn from a one-site conditional under some $\mu_{i}(\cdot\mid T_{<t})$, hence Section 4.1 applies uniformly over the conditioning. Therefore, for each queried vertex, the probability of landing in the relevant boundary strip is at most $C_{\mathrm{bdry}}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}\cdot\exp\!\left(-c_{\mathrm{bdry}}\cdot\delta\cdot\frac{(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2}\right)$. A union bound over the at most $q\leq Q$ queried vertices gives the stated bound on $\eta_{\partial,\mathcal{A}}$.
Finally, choosing $\gamma=C\sqrt{\log(100Q^{2})}$ with $C$ large enough makes the RHS at most $0.01$. ∎
4.4 Prefix-conditioned leakage via strong convexity (Green-function comparison)
This subsection bounds adaptive information leakage by comparing one-site marginals under different hidden indices via strong convexity.
For each $i$ and site $v$, denote the one-site conditional kernel by
$K_{v}^{(i)}(\cdot\mid x_{-v}).$
We will not use Dobrushin uniqueness. Instead, we compare one-site marginals under transcript prefixes using only the strong convexity from the quadratic part $\frac{\alpha}{2}\sum_{v}f(v)^{2}+\lambda\sum_{(\ell,r)\in E}(f(\ell)-f(r))^{2}$.
Let $\Gamma=(L\cup R,E)$ be the underlying undirected graph and let $\mathbf{L}$ be its (combinatorial) Laplacian matrix: $\mathbf{L}_{v,v}=\deg(v)$, $\mathbf{L}_{v,u}=-1$ if $\{v,u\}\in E$, and 0 otherwise. For a partial transcript condition $\tau$ fixing a set $S(\tau)\subseteq L\cup R$ of coordinates, let $V_{\tau}:=(L\cup R)\setminus S(\tau)$ be the set of unfixed sites and let $\mathbf{L}_{\tau}$ be the principal submatrix of $\mathbf{L}$ indexed by $V_{\tau}$. Define
$\mathbf{M}_{\tau}:=\alpha I_{V_{\tau}}+2\lambda\mathbf{L}_{\tau},\qquad G_{\tau}:=\mathbf{M}_{\tau}^{-1}.$
Row-sum bound for $G_{\tau}$. The matrix $\mathbf{M}_{\tau}$ has positive diagonal and nonpositive off-diagonal entries, and it is strictly diagonally dominant because $\alpha>0$; hence it is a nonsingular $M$-matrix and $G_{\tau}$ is entrywise nonnegative. Moreover, writing $\mathbf{1}$ for the all-ones vector on $V_{\tau}$, we have $\mathbf{L}_{\tau}\mathbf{1}\geq 0$ entrywise (edges to fixed vertices contribute only to the diagonal), and therefore $\mathbf{M}_{\tau}\mathbf{1}=\alpha\mathbf{1}+2\lambda\mathbf{L}_{\tau}\mathbf{1}\geq\alpha\mathbf{1}$. Multiplying by the entrywise nonnegative $G_{\tau}$ yields the coordinatewise bound
$G_{\tau}\mathbf{1}\leq\frac{1}{\alpha}\mathbf{1},\qquad\text{i.e.,}\qquad\sum_{u\in V_{\tau}}(G_{\tau})_{v,u}\leq\frac{1}{\alpha}\ \ \ (\forall v\in V_{\tau}).$ (22)
For $i,j\in[s]$, define local KL discrepancies
$b_{u}^{(i,j)}:=\sup_{x_{-u}\in[-B,B]^{V\setminus\{u\}}}\mathrm{KL}\!\big(K_{u}^{(i)}(\cdot\mid x_{-u})\,\|\,K_{u}^{(j)}(\cdot\mid x_{-u})\big).$
Lemma 4.13 (Local one-site discrepancy).
For every $u$ and $i,j$,
$b_{u}^{(i,j)}\leq C_{b}\frac{\beta^{2}(B+\gamma)^{2}}{m_{u}},$
for an absolute constant $C_{b}>0$.
Proof.
Fix $u,i,j,x_{-u}$. Write the one-site conditionals as densities on [B,B][-B,B]:
$p_{i}(x)\propto e^{-h_{i}(x)},\qquad p_{j}(x)\propto e^{-h_{j}(x)}.$
Then
$\partial_{x}\log p_{i}(x)-\partial_{x}\log p_{j}(x)=-(h_{i}^{\prime}(x)-h_{j}^{\prime}(x)).$
The two one-site log-densities differ only in terms corresponding to incident edges whose matching type changes from $i$ to $j$; this is at most two terms. For one such changed edge with neighbor value $y\in[-B,B]$, the hinge derivative contribution is of the form $\pm 2\beta(x-y+\gamma)_{+}$ or $\pm 2\beta(\gamma-(x-y))_{+}$, so for $x\in[-B,B]$ its magnitude is at most $2\beta(2B+\gamma)\leq C\beta(B+\gamma)$. Summing over at most two changed edges gives
$\sup_{x\in[-B,B]}\big|h_{i}^{\prime}(x)-h_{j}^{\prime}(x)\big|\leq C_{0}\beta(B+\gamma),$
hence
$\sup_{x\in[-B,B]}\big|\partial_{x}\log p_{i}(x)-\partial_{x}\log p_{j}(x)\big|\leq C_{0}\beta(B+\gamma).$
Each one-site conditional is $m_{u}$-strongly log-concave (Section 4.1), so the one-dimensional entropy estimate gives
$\mathrm{KL}(p_{i}\|p_{j})\leq\frac{1}{2m_{u}}\,\sup_{x}\big|\partial_{x}\log p_{i}(x)-\partial_{x}\log p_{j}(x)\big|^{2}\leq C_{b}\frac{\beta^{2}(B+\gamma)^{2}}{m_{u}}.$
Taking the supremum over $x_{-u}$ yields the claim. ∎
To apply the Green-function/interpolation step below, we first work with smooth ($C^{2}$) approximations: the original model has hinge kinks and a hard box truncation, so direct differentiation-based identities are not available in full regularity. We then pass to the original model by letting the smoothing parameter tend to 0.
Lemma 4.14 (Uniform stability of one-site KL under smooth approximation).
Let $B>0$ and let $\Theta$ be a compact parameter set. For $k\in\{1,2\}$ let $h^{(k)}:[-B,B]\times\Theta\to\mathbb{R}$ be continuous and let $h_{\varepsilon}^{(k)}:[-B,B]\times\Theta\to\mathbb{R}$ be continuous such that $h_{\varepsilon}^{(k)}\to h^{(k)}$ uniformly on $B,B]\times\Theta$ as $\varepsilon\downarrow 0$. Define the associated normalized densities on [B,B][-B,B] by
$K_{\varepsilon}^{(k)}(x\mid\theta):=\frac{\exp(-h_{\varepsilon}^{(k)}(x,\theta))}{\int_{-B}^{B}\exp(-h_{\varepsilon}^{(k)}(y,\theta))\,dy},\qquad K^{(k)}(x\mid\theta):=\frac{\exp(-h^{(k)}(x,\theta))}{\int_{-B}^{B}\exp(-h^{(k)}(y,\theta))\,dy}.$
Then
$\sup_{\theta\in\Theta}\Big|\operatorname{KL}(K_{\varepsilon}^{(1)}(\cdot\mid\theta)\,\|\,K_{\varepsilon}^{(2)}(\cdot\mid\theta))-\operatorname{KL}(K^{(1)}(\cdot\mid\theta)\,\|\,K^{(2)}(\cdot\mid\theta))\Big|\to 0\qquad(\varepsilon\downarrow 0),$
and hence
$\lim_{\varepsilon\downarrow 0}\sup_{\theta\in\Theta}\operatorname{KL}(K_{\varepsilon}^{(1)}(\cdot\mid\theta)\,\|\,K_{\varepsilon}^{(2)}(\cdot\mid\theta))=\sup_{\theta\in\Theta}\operatorname{KL}(K^{(1)}(\cdot\mid\theta)\,\|\,K^{(2)}(\cdot\mid\theta)).$
Proof.
Uniform convergence of $h_{\varepsilon}^{(k)}$ implies uniform convergence of the unnormalized densities $\exp(-h_{\varepsilon}^{(k)})$ and of the partition functions $Z_{\varepsilon}^{(k)}(\theta)=\int_{-B}^{B}\exp(-h_{\varepsilon}^{(k)}(y,\theta))dy$ on the compact set $\Theta$. Therefore $\log K_{\varepsilon}^{(k)}(x\mid\theta)\to\log K^{(k)}(x\mid\theta)$ uniformly on $B,B]\times\Theta$, and in particular the log-likelihood ratios $\log\frac{K_{\varepsilon}^{(1)}}{K_{\varepsilon}^{(2)}}$ converge uniformly and are uniformly bounded in absolute value. Since $K_{\varepsilon}^{(1)}(\cdot\mid\theta)$ are probability densities, the KL integrands $K_{\varepsilon}^{(1)}\log\frac{K_{\varepsilon}^{(1)}}{K_{\varepsilon}^{(2)}}$ converge uniformly and are dominated by an integrable constant, uniformly over $\theta$. Dominated convergence yields the stated uniform convergence of KL. ∎
Lemma 4.15 (Comparison under partial conditioning).
For every partial transcript condition $\tau$ (fixing any subset of coordinates), every unqueried site $v\in V_{\tau}$, and every indices $i,j$,
$\mathrm{KL}\!\left(\mathcal{L}_{\mu_{i}}(f(v)\mid\tau)\,\middle\|\,\mathcal{L}_{\mu_{j}}(f(v)\mid\tau)\right)\leq\sum_{u\in V_{\tau}}(G_{\tau})_{v,u}\,b_{u}^{(i,j)}.$
Proof.
Fix $i,j\in[s]$, a transcript $\tau$, and a site $v\in V_{\tau}$. Let $\nu_{k}:=\mu_{k}(\cdot\mid\tau)$ be the conditional Gibbs measure on the remaining coordinates $V_{\tau}$, for $k\in\{i,j\}$. Its density has the form
$\nu_{k}(dx)\ =\ \frac{1}{Z_{k}}\,\exp\!\big(-H_{k}^{\tau}(x)\big)\,\mathbf{1}\{x\in[-B,B]^{V_{\tau}}\}\,dx,$
where $H_{k}^{\tau}$ is the conditional Hamiltonian obtained from $H_{k}$ by plugging in the fixed values from $\tau$. Because every edge-term (quadratic smoothing and hinge-square) is convex in the spins, both $H_{i}^{\tau}$ and $H_{j}^{\tau}$ are convex, and moreover their Hessians (in the sense of distributions) satisfy the same uniform lower bound
$\nabla^{2}H_{k}^{\tau}(x)\ \succeq\ \mathbf{M}_{\tau}\ :=\ \alpha I+2\lambda L_{\tau},\qquad k\in\{i,j\},$
for all $x\in(-B,B)^{V_{\tau}}$. In particular, $\mathbf{G}_{\tau}=\mathbf{M}_{\tau}^{-1}$ is well-defined.
Let $K_{u}^{(k)}(\cdot\mid x_{-u})$ be the one-site conditional kernel at $u$ under $\nu_{k}$. Define
$b_{u}^{(i,j)}\ :=\ \sup_{x_{-u}}\operatorname{KL}\!\left(K_{u}^{(i)}(\cdot\mid x_{-u})\,\big\|\,K_{u}^{(j)}(\cdot\mid x_{-u})\right).$
The desired inequality
$\operatorname{KL}\!\left(\nu_{i}(X_{v})\,\big\|\,\nu_{j}(X_{v})\right)\ \leq\ \sum_{u\in V_{\tau}}(\mathbf{G}_{\tau})_{v,u}\,b_{u}^{(i,j)}$
is proved in the literature as an entropy influence / Green–function comparison inequality for uniformly convex Gibbs measures. Since this lemma is used as an analytic input for the leakage bound, we briefly justify it here and also explain why the hinge kinks and hard box truncation do not invalidate the argument.
Step 1 (smooth approximation of kinks and truncation). Fix $\varepsilon>0$. Replace each hinge-square term $t\mapsto(t)_{+}^{2}$ by a convex $C^{2}$ mollification $t\mapsto\psi_{\varepsilon}(t)$ which agrees with $(t)_{+}^{2}$ outside an $\varepsilon$-neighborhood of the kink at $t=0$. Also replace the hard indicator $\mathbf{1}\{x\in[-B,B]\}$ by adding a convex $C^{2}$ barrier potential $\Psi_{\varepsilon}$ that is 0 on [B,B][-B,B] and diverges to $\infty$ as $|x|\uparrow B$. Let $H_{k,\varepsilon}^{\tau}$ be the resulting $C^{2}$ Hamiltonian on $\mathbb{R}^{V_{\tau}}$ and let $\nu_{k,\varepsilon}$ be the corresponding smooth Gibbs measure (without an indicator). Because we only replace terms by more convex ones, we preserve the same uniform curvature lower bound
$\nabla^{2}H_{k,\varepsilon}^{\tau}(x)\ \succeq\ \mathbf{M}_{\tau}\qquad(\forall x,\ k\in\{i,j\}).$
Moreover $\nu_{k,\varepsilon}\Rightarrow\nu_{k}$ as $\varepsilon\downarrow 0$ (the barrier localizes to the box and the mollified hinges converge pointwise). In particular, for every fixed boundary condition $x_{-u}$, the one-site conditional densities $K_{u,\varepsilon}^{(k)}(\cdot\mid x_{-u})$ converge in $L^{1}([-B,B])$ to $K_{u}^{(k)}(\cdot\mid x_{-u})$. In Step 3 we record a uniformity statement that allows us to pass the supremum over $x_{-u}$.
Step 2 (Green–function comparison in the smooth uniformly convex case). For the smooth measures $\nu_{k,\varepsilon}$ the inequality
$\operatorname{KL}\!\left(\nu_{i,\varepsilon}(X_{v})\,\big\|\,\nu_{j,\varepsilon}(X_{v})\right)\ \leq\ \sum_{u\in V_{\tau}}(\mathbf{G}_{\tau})_{v,u}\,b_{u,\varepsilon}^{(i,j)}$
follows from the standard interpolation argument: interpolate $H_{t,\varepsilon}^{\tau}:=(1-t)H_{i,\varepsilon}^{\tau}+tH_{j,\varepsilon}^{\tau}$ and differentiate the $v$-marginal likelihood ratio along this path. The derivative can be expressed as a covariance between a $v$-measurable quantity and a local perturbation affecting only one site $u$. Applying the Brascamp–Lieb covariance inequality (equivalently, the Helffer–Sjöstrand representation) under the uniform bound $\nabla^{2}H_{t,\varepsilon}^{\tau}\succeq\mathbf{M}_{\tau}$ shows that the sensitivity of the $v$-marginal to such a perturbation at $u$ is controlled by the Green function entry $(\mathbf{G}_{\tau})_{v,u}$. Integrating over $t\in[0,1]$ and using that the entropy cost of swapping the $u$-one-site kernels is at most $b_{u,\varepsilon}^{(i,j)}$ yields the displayed bound.
Step 3 (remove the smoothing). Finally, let $\varepsilon\downarrow 0$. The barrier potential $\Psi_{\varepsilon}$ forces $\nu_{k,\varepsilon}$ to concentrate on $B,B]^{V_{\tau}}$, while the mollified hinges converge pointwise to the original hinge-square terms on this compact domain. In particular, $\nu_{k,\varepsilon}\Rightarrow\nu_{k}$ weakly for $k\in\{i,j\}$, and the corresponding one-site conditional densities $K_{u,\varepsilon}^{(k)}(\cdot\mid x_{-u})$ converge pointwise on [B,B][-B,B] to $K_{u}^{(k)}(\cdot\mid x_{-u})$ for every fixed $x_{-u}$.
Moreover, because $u$ interacts with only finitely many neighbors, the boundary condition $x_{-u}$ enters each one-site potential only through the neighbor vector $\theta:=x_{\mathrm{bdry}_{u}}\in\Theta:=[-B,B]^{\deg(u)}$. Our mollification $\psi_{\varepsilon}$ and barrier $\Psi_{\varepsilon}$ can be chosen so that, for each $k\in\{i,j\}$, the one-site conditional log-densities $h_{u,\varepsilon}^{(k)}(x\mid\theta)$ converge uniformly to $h_{u}^{(k)}(x\mid\theta)$ on $B,B]\times\Theta$. Therefore Section 4.4 applies and yields uniform convergence of the one-site KL cost as a function of $\theta$, which in particular justifies exchanging $\lim$ and $\sup$:
$\lim_{\varepsilon\downarrow 0}b_{u,\varepsilon}^{(i,j)}\;=\;b_{u}^{(i,j)}.$
Finally, KL is lower semicontinuous under weak convergence of the corresponding marginals, so passing to the limit in Step 2 gives the desired inequality for $\nu_{i},\nu_{j}$. ∎
Definition 4.16 (Per-query leakage scale).
$\kappa_{Q}:=\frac{C_{b}}{\alpha}\cdot\frac{\beta^{2}(B+\gamma)^{2}}{m_{\min}}.$
Lemma 4.17 (Leakage bound up to budget $Q$).
For every deterministic adaptive strategy making $q\leq Q$ queries,
$I(I;T)\leq q\,\kappa_{Q}.$
Proof.
Fix $t\leq q$ and condition on $T_{<t}=\tau$. Let $v_{t}(\tau)$ be the queried vertex and $P_{i}^{\tau}:=\mathcal{L}(\mathrm{Ans}_{t}\mid I=i,T_{<t}=\tau)$. Then
$I(I;\mathrm{Ans}_{t}\mid T_{<t}=\tau)\leq\max_{i,j}\mathrm{KL}(P_{i}^{\tau}\|P_{j}^{\tau}).$
By Section 4.4 and Section 4.4,
$\mathrm{KL}(P_{i}^{\tau}\|P_{j}^{\tau})\leq\sum_{u\in V_{\tau}}(G_{\tau})_{v_{t}(\tau),u}\,b_{u}^{(i,j)}\leq\left(\max_{u\in V_{\tau}}b_{u}^{(i,j)}\right)\sum_{u\in V_{\tau}}(G_{\tau})_{v_{t}(\tau),u}.$
Using Section 4.4, $m_{u}\geq m_{\min}$, and the row-sum bound (22), we get
$\mathrm{KL}(P_{i}^{\tau}\|P_{j}^{\tau})\leq C_{b}\frac{\beta^{2}(B+\gamma)^{2}}{m_{\min}}\cdot\frac{1}{\alpha}=\kappa_{Q}.$
So each summand in
$I(I;T)=\sum_{t=1}^{q}I(I;\mathrm{Ans}_{t}\mid T_{<t})$
is at most $\kappa_{Q}$, hence $I(I;T)\leq q\kappa_{Q}$. ∎
4.5 Unique decoding and standard Fano
This subsection turns witnessed violated edges into an index decoder and then applies Fano’s inequality to lower-bound the required information.
Because the matchings are edge-disjoint, every $e\in\bigcup_{i=1}^{s}M_{i}$ has a unique index $J(e)\in[s]$ with $e\in M_{J(e)}$. Define
$\mathsf{SpurQry}:=\{\exists\text{ queried violated edge in }E\setminus M_{I}\}.$
Let
$\eta_{\partial}:=\Pr[\mathsf{BdryQry}].$
Definition 4.18 (Transcript decoder).
Given transcript $T$, define $\widehat{I}(T)$:
  1. 1.
    if $T$ contains at least one witnessed violated queried edge, take the first such edge $e$ under a fixed rule; if $e\in\bigcup_{i}M_{i}$, output $J(e)$, otherwise output $1$;
  2. 2.
    if no witnessed violated queried edge exists, output $1$.
Lemma 4.19 (Worst-case to distributional rejection).
Let $\mathcal{A}$ be a (possibly randomized) one-sided $\varepsilon$-tester with worst-case soundness $2/3$ on $\varepsilon$-far inputs. If a distribution $\mathcal{D}$ over inputs satisfies $\Pr_{f\sim\mathcal{D}}[f\text{ is }\varepsilon\text{-far}]\geq 1-\delta$, then
$\Pr_{f\sim\mathcal{D},\rho}[\mathcal{A}^{\rho}\ \mathrm{rejects}\ f]\geq(1-\delta)\cdot\frac{2}{3},$
where $\rho$ is the internal randomness of $\mathcal{A}$.
Proof.
Condition on the event $\{f\text{ is }\varepsilon\text{-far}\}$. On this event, worst-case soundness gives rejection probability at least $2/3$. Multiply by its probability at least $1-\delta$. ∎
Lemma 4.20 (Soundness implies decoding success).
For every deterministic adaptive one-sided tester,
$\Pr[\widehat{I}(T)=I]\geq\Pr[\mathrm{reject}]-\Pr[\mathsf{SpurQry}].$
Hence, if soundness is $2/3$ on $\varepsilon$-far inputs ($\varepsilon=\varepsilon_{0}/200$), and if Section 4.3 and Section 4.3 apply, then
$\Pr[\widehat{I}(T)=I]\geq 0.65-\eta_{\partial}.$
Proof.
On $\{\mathrm{reject}\}\setminus\mathsf{SpurQry}$, every witnessed violated queried edge belongs to $M_{I}$. By one-sidedness (Section 4), rejection implies existence of a witnessed violated queried edge, so the decoder outputs $I$. Hence
$\{\mathrm{reject}\}\setminus\mathsf{SpurQry}\subseteq\{\widehat{I}(T)=I\},$
which gives the first inequality. By Section 4.3, under the hard prior we have $\Pr[\varepsilon\text{-far}]\geq 0.99$. Section 4.5 gives $\Pr[\mathrm{reject}]\geq 0.99\cdot(2/3)>0.66$. Section 4.3 gives $\Pr[\mathsf{SpurQry}]\leq\eta_{\partial}+0.01$. ∎
Lemma 4.21 (Standard Fano (uniform prior)).
Let $I\sim\mathrm{Unif}([s])$, transcript $T$, and decoder $\widehat{I}=\widehat{I}(T)$. If $P_{e}:=\Pr[\widehat{I}\neq I]$, then
$I(I;T)\geq(1-P_{e})\log s-1.$
Proof.
Apply Fano’s inequality to the decoder $\widehat{I}(T)$ for the uniform prior $I\sim\mathrm{Unif}([s])$. With error probability $P_{e}=\Pr[\widehat{I}\neq I]$, this gives
$H(I\mid T)\leq h_{2}(P_{e})+P_{e}\log(s-1)\leq 1+P_{e}\log s,$
where $h_{2}(\cdot)$ is the binary entropy function (and $h_{2}(P_{e})\leq 1$). Since $H(I)=\log s$,
$I(I;T)=H(I)-H(I\mid T)\geq(1-P_{e})\log s-1.$
4.6 Budgeted one-sided lower bound
We now combine the bounds controlling spurious queried violations and leakage from the previous subsections to obtain a fixed-budget lower bound for adaptive one-sided testers.
Theorem 4.22 (Fixed-budget information lower bound).
Fix $Q\geq 1$ and parameters $(\alpha,\lambda,\beta,\gamma,B)$ as above. Assume the hypotheses needed for Section 4.2 and Section 4.4 (in particular (15) and (16)), and assume
$p_{\mathrm{match}}\leq 10^{-4},\qquad p_{\mathrm{spur}}\leq\frac{1}{100Q^{2}}.$
Let $\mathcal{A}$ be any deterministic adaptive one-sided strategy making $q\leq Q$ queries, and set
$\rho_{\mathcal{A}}:=\Pr_{\mathcal{A}}[\mathrm{reject}],\qquad\eta_{\partial,\mathcal{A}}:=\Pr_{\mathcal{A}}[\mathsf{BdryQry}].$
Then
$q\kappa_{Q}\geq(\rho_{\mathcal{A}}-\eta_{\partial,\mathcal{A}}-0.01)\log s-1.$
In particular, with $\varepsilon:=\varepsilon_{0}/200$, if $\mathcal{A}$ is an $\varepsilon$-tester with soundness $2/3$ on $\varepsilon$-far inputs, then
$q\kappa_{Q}\geq(0.65-\eta_{\partial,\mathcal{A}})\log s-1.$
Hence, if every deterministic adaptive one-sided $\varepsilon$-tester with at most $Q$ queries satisfies $\eta_{\partial,\mathcal{A}}\leq\eta_{0}$ and
$Q\kappa_{Q}<(0.65-\eta_{0})\log s-1,$
then no such tester with at most $Q$ queries exists.
Proof.
By Section 4.5,
$\Pr[\widehat{I}(T)=I]\geq\rho_{\mathcal{A}}-\Pr[\mathsf{SpurQry}].$
Section 4.3 gives $\Pr[\mathsf{SpurQry}]\leq\eta_{\partial,\mathcal{A}}+0.01$, so
$\Pr[\widehat{I}(T)=I]\geq\rho_{\mathcal{A}}-\eta_{\partial,\mathcal{A}}-0.01.$
Thus
$P_{e}:=\Pr[\widehat{I}(T)\neq I]\leq 1-\rho_{\mathcal{A}}+\eta_{\partial,\mathcal{A}}+0.01.$
Applying Section 4.5,
$I(I;T)\geq(1-P_{e})\log s-1\geq(\rho_{\mathcal{A}}-\eta_{\partial,\mathcal{A}}-0.01)\log s-1.$
Section 4.4 gives $I(I;T)\leq q\kappa_{Q}$, proving the first claim.
If $\mathcal{A}$ is a sound $\varepsilon$-tester, then Section 4.3 implies $\Pr[\varepsilon\text{-far}]\geq 0.99$, and Section 4.5 gives $\rho_{\mathcal{A}}\geq 0.99\cdot(2/3)>0.66$. Substitute this into the first bound to get $q\kappa_{Q}\geq(0.65-\eta_{\partial,\mathcal{A}})\log s-1$. ∎
4.7 An $\Omega(\sqrt{n})$ instantiation and the mean-control bound
To apply the fixed-budget lower bound Theorem 4.22, we verify the admissible-parameter conditions and the required smallness bounds for $(p_{\mathrm{match}},p_{\mathrm{spur}},p_{\mathrm{mean}},\eta_{0})$. In this section we give an explicit dense $C_{4}$-free PMRS base graph and verify all of these quantities. To bound the mean-control term $p_{\mathrm{mean}}$, we combine the Lipschitz and concentration bounds in Section 4.7 with the additional centering estimate (Section 4.7) available in this dense regime. This yields the $\Omega(\sqrt{n})$ adaptive one-sided lower bound.
$C_{4}$-free PMRS base graphs from difference-free shifts.
Recall the PMRS construction in Section 3.4 (“shift matchings”) which produces a family of positive matchings indexed by a set $A$ of integer vectors. For this shift-matching construction, the resulting bipartite graph is $C_{4}$-free provided $A$ is difference-free.
Lemma 4.23 (Difference-free shift sets yield $C_{4}$-free PMRS graphs).
Consider the “shift-matching” base graph of Section 3.4 built from a set of shifts $A\subseteq\mathbb{Z}^{k}$ and let $U$ be the union graph of the matchings $\{M_{\mathbf{a}}\}_{\mathbf{a}\in A}$. If $A$ is difference-free in the sense that
$(A-A)\cap A\ \subseteq\ \{0\},$
then the underlying undirected graph of $U$ has no $4$-cycles. Moreover, each $M_{\mathbf{a}}$ is a positive matching in $U$ (as in Section 3.4).
Proof.
A $4$-cycle in the underlying undirected graph has the form
[…]
where the labels $\mathbf{a},\mathbf{b}\in A$ indicate which shift-matching each traversed edge belongs to. (Traversing an edge backwards simply swaps endpoints but keeps the same shift label.) Write
$\ell=(x,z)_{L},\qquad\ell^{\prime}=(x^{\prime},z^{\prime})_{L}.$
By the definition of shift matchings, an edge of label $t$ maps $(u,w)_{L}$ to $(u+t,\;w+\|t\|_{2}^{2})_{R}$. Hence
$r=(x+\mathbf{b},\;z+\|\mathbf{b}\|_{2}^{2})_{R}=(x^{\prime}+\mathbf{a},\;z^{\prime}+\|\mathbf{a}\|_{2}^{2})_{R},$
and
$r^{\prime}=(x+\mathbf{a},\;z+\|\mathbf{a}\|_{2}^{2})_{R}=(x^{\prime}+\mathbf{b},\;z^{\prime}+\|\mathbf{b}\|_{2}^{2})_{R}.$
Comparing first coordinates gives
$x^{\prime}-x=\mathbf{b}-\mathbf{a},\qquad x^{\prime}-x=\mathbf{a}-\mathbf{b}.$
Therefore $\mathbf{a}=\mathbf{b}$. But then both edges incident to $\ell$ in the cycle, namely $\ell r$ and $\ell r^{\prime}$, belong to the same matching $M_{\mathbf{a}}$, impossible because $M_{\mathbf{a}}$ is a matching. So no $4$-cycle can exist.
Positivity of each $M_{\mathbf{a}}$ follows from the same weight witness as in Section 3.4: for an edge from matching $M_{\mathbf{b}}$ one has $w_{\mathbf{a}}(u)+w_{\mathbf{a}}(v)=\tfrac{1}{2}-\|\mathbf{b}-\mathbf{a}\|_{2}^{2}$, which is positive iff $\mathbf{b}=\mathbf{a}$. ∎
A dense explicit difference-free set in $k=2$.
Fix integers $N\geq 1$ and $P\geq 1$ with $P\leq N/10$. Define the set of shifts
$A\;:=\;\bigl\{\mathbf{a}=(a_{1},a_{2})\in\{0,1,\dots,P\}^{2}\setminus\{(0,0)\}:\ a_{1}>\lfloor P/2\rfloor\bigr\},\qquad s:=|A|=\Theta(P^{2}).$
Then $A$ is difference-free: if $\mathbf{a}\neq\mathbf{b}$ are in $A$, then $(\mathbf{a}-\mathbf{b})_{1}\in[-\lfloor P/2\rfloor,\lfloor P/2\rfloor]$, so $\mathbf{a}-\mathbf{b}\notin A$ and $\mathbf{b}-\mathbf{a}\notin A$. By Section 4.7, the associated shift-matching union graph is $C_{4}$-free.
The explicit base graph and trimming.
We specialize the explicit $k=2$ shift-matching construction from Theorem 3.9. Let
$L=[N]^{2}\times[N^{2}],\qquad R=[N]^{2}\times[N^{2}],$
and for each $\mathbf{a}\in A$ define the matching
$M_{\mathbf{a}}:=\bigl\{\,((x,z)_{L},(x+\mathbf{a},\,z+\|\mathbf{a}\|_{2}^{2})_{R}):\ x,x+\mathbf{a}\in[N]^{2},\ z,z+\|\mathbf{a}\|_{2}^{2}\in[N^{2}]\,\bigr\}.$
Let $U=(L,R;E)$ with $E:=\bigcup_{\mathbf{a}\in A}M_{\mathbf{a}}$.
We extract a balanced dense subgraph by pruning by degree and a simple balancing step. A simultaneously perfect/regular subgraph need not exist for multiple distinct shifts, and this subgraph is sufficient for the lower-bound parameters.
Lemma 4.24 (Balanced dense subgraph of the $k=2$ shift-matching base graph).
Fix $N,P\in\mathbb{N}$ with $P\leq N/10$, and let $A\subseteq\{0,1,\dots,P\}^{2}\setminus\{(0,0)\}$ be difference-free, with $s:=|A|$. Let $L=R=[N]^{2}\times[N^{2}]$, and let $U=(L,R;E)$ be the union graph $E=\bigcup_{\mathbf{a}\in A}M_{\mathbf{a}}$ of the $k=2$ shift matchings from Theorem 3.9.
Then there exist subsets $L^{\circ}\subseteq L$, $R^{\circ}\subseteq R$ with $|L^{\circ}|=|R^{\circ}|=:n_{0}$ and a subset $A^{\circ}\subseteq A$ with $|A^{\circ}|\geq\frac{c_{\delta}}{2-c_{\delta}}s$ such that, writing $U^{\circ}=(L^{\circ},R^{\circ};E^{\circ})$ for the induced subgraph and $M_{\mathbf{a}}^{\circ}:=M_{\mathbf{a}}\cap E^{\circ}$ for $\mathbf{a}\in A^{\circ}$, we have:
  1. 1.
    $U^{\circ}$ is $C_{4}$-free.
  2. 2.
    For every $\mathbf{a}\in A^{\circ}$, $M_{\mathbf{a}}^{\circ}$ is a positive matching in $U^{\circ}$.
  3. 3.
    Each $M_{\mathbf{a}}^{\circ}$ has size at least $\varepsilon_{0}n_{0}$ for an absolute constant $\varepsilon_{0}>0$.
  4. 4.
    Degrees are bounded as $c_{\delta}s\leq\deg_{U^{\circ}}(v)\leq s$ for all vertices $v$ (thus $\delta=\Omega(s)$ and $\Delta\leq s$), for an absolute constant $c_{\delta}>0$.
In particular, under the dense choice $P=\lfloor N/10\rfloor$ we have $s=\Theta(N^{2})$ and $n_{0}=\Theta(N^{4})$, hence $s=\Theta(\sqrt{n_{0}})$ and $\delta=\Theta(s)$.
Proof.
For each $\mathbf{a}=(a_{1},a_{2})\in A$, the matching $M_{\mathbf{a}}$ consists of all edges
$(x,z)\mapsto(x+\mathbf{a},\;z+\|\mathbf{a}\|_{2}^{2})$
with $x,x+\mathbf{a}\in[N]^{2}$ and $z,z+\|\mathbf{a}\|_{2}^{2}\in[N^{2}]$. Hence
$|M_{\mathbf{a}}|=(N-a_{1})(N-a_{2})\,(N^{2}-\|\mathbf{a}\|_{2}^{2})\ \geq\ (N-P)^{2}\,(N^{2}-2P^{2}).$
Let $c_{0}:=(1-P/N)^{2}(1-2P^{2}/N^{2})$. For $P\leq N/10$ we have $c_{0}\geq 0.79$. Since the matchings are edge-disjoint, the union has
$|E|=\sum_{\mathbf{a}\in A}|M_{\mathbf{a}}|\ \geq\ c_{0}\,s\,N^{4}.$
Because $|L|=|R|=N^{4}$, the average degree on each side is at least $c_{0}s$.
Now perform the standard pruning-by-degree procedure: starting from $U$, iteratively delete any vertex (on either side) whose current degree is less than $(c_{0}/4)s$, together with all incident edges, until no such vertex remains. Let the resulting subgraph be $U_{1}=(L_{1},R_{1};E_{1})$. By construction, every vertex in $U_{1}$ has degree at least $(c_{0}/4)s$, and trivially every degree is at most $s$. Each deletion step removes fewer than $(c_{0}/4)s$ edges, so after at most $|L|+|R|=2N^{4}$ steps we have
$|E_{1}|\ \geq\ |E|-2N^{4}\cdot\frac{c_{0}}{4}s\ \geq\ \frac{c_{0}}{2}\,s\,N^{4}.$
Since each left vertex has degree at most $s$, we also have $|E_{1}|\leq s\,\min\{|L_{1}|,|R_{1}|\}$, and therefore
$n_{1}:=\min\{|L_{1}|,|R_{1}|\}\ \geq\ \frac{|E_{1}|}{s}\ \geq\ \frac{c_{0}}{2}N^{4}.$
Moreover, because $U_{1}$ has minimum degree at least $(c_{0}/4)s$ and maximum degree at most $s$, the ratio of side sizes is bounded by
$\frac{\max\{|L_{1}|,|R_{1}|\}}{\min\{|L_{1}|,|R_{1}|\}}\ \leq\ \frac{s}{(c_{0}/4)s}=\frac{4}{c_{0}}.$
To balance the sides, let $n_{0}:=n_{1}$ and (wlog) assume $|L_{1}|\geq|R_{1}|=n_{0}$. Choose a subset $L^{\circ}\subseteq L_{1}$ of size $n_{0}$ uniformly at random, and set $R^{\circ}:=R_{1}$. For any fixed $r\in R^{\circ}$, the number of its neighbors retained in $L^{\circ}$ is hypergeometric with mean
$\mathbb{E}[\deg_{U^{\circ}}(r)]\geq\deg_{U_{1}}(r)\cdot\frac{|L^{\circ}|}{|L_{1}|}\geq\frac{c_{0}}{4}s\cdot\frac{n_{0}}{|L_{1}|}\geq\frac{c_{0}}{4}s\cdot\frac{1}{4/c_{0}}=\frac{c_{0}^{2}}{16}s.$
A standard hypergeometric Chernoff bound implies that
$\Pr\!\left[\deg_{U^{\circ}}(r)<\frac{c_{0}^{2}}{32}s\right]\leq\exp(-\Omega(s)).$
Taking a union bound over all $r\in R^{\circ}$ shows that with positive probability every right vertex has degree at least $(c_{0}^{2}/32)s$ in $U^{\circ}$. Left degrees are unchanged when we subsample $L_{1}$ (we did not delete any right vertices), so every left vertex in $L^{\circ}$ still has degree at least $(c_{0}/4)s$. Thus, in the balanced graph $U^{\circ}$ we have
$\deg_{U^{\circ}}(v)\ \geq\ c_{\delta}s\qquad\text{with}\qquad c_{\delta}:=\frac{c_{0}^{2}}{32},$
and also $\deg_{U^{\circ}}(v)\leq s$, i.e., $\Delta\leq s$.
Since $A$ is difference-free, the original union graph is $C_{4}$-free by Section 4.7, and this property is preserved under taking subgraphs, giving (1). Positivity of each $M_{\mathbf{a}}$ in $U$ comes from Theorem 3.9, and it is preserved under restricting to a subgraph, giving (2).
Finally, the restricted matchings $\{M_{\mathbf{a}}^{\circ}\}_{\mathbf{a}\in A}$ remain edge-disjoint and satisfy
$\sum_{\mathbf{a}\in A}|M_{\mathbf{a}}^{\circ}|=|E^{\circ}|\geq|L^{\circ}|\cdot\min_{v\in L^{\circ}}\deg_{U^{\circ}}(v)\geq c_{\delta}sn_{0}.$
Hence the average of $|M_{\mathbf{a}}^{\circ}|$ over $\mathbf{a}\in A$ is at least $c_{\delta}n_{0}$. Let $A^{\circ}\subseteq A$ be the set of shifts with $|M_{\mathbf{a}}^{\circ}|\geq(c_{\delta}/2)n_{0}$. Since always $|M_{\mathbf{a}}^{\circ}|\leq n_{0}$, the averaging bound implies a  constant fraction of shifts are dense. Indeed, writing $p:=|A^{\circ}|/s$ and using $|M_{\mathbf{a}}^{\circ}|\leq n_{0}$ for $\mathbf{a}\in A^{\circ}$ and $|M_{\mathbf{a}}^{\circ}|<(c_{\delta}/2)n_{0}$ for $\mathbf{a}\notin A^{\circ}$, we have
$\frac{1}{s}\sum_{\mathbf{a}\in A}|M_{\mathbf{a}}^{\circ}|\leq p\cdot n_{0}+(1-p)\cdot(c_{\delta}/2)n_{0}.$
Since the left-hand side is at least $c_{\delta}n_{0}$, rearranging gives
$p\geq\frac{c_{\delta}}{2-c_{\delta}}.$
Thus $|A^{\circ}|\geq\frac{c_{\delta}}{2-c_{\delta}}\,s$, and each $M_{\mathbf{a}}^{\circ}$ has size at least $\varepsilon_{0}n_{0}$ with $\varepsilon_{0}:=c_{\delta}/2$. This proves (3)–(4) and completes the proof. ∎
In what follows, we fix the dense choice $P=\lfloor N/10\rfloor$ and work with a balanced core $U^{\circ}$ as in Section 4.7. We write $n_{0}:=|L^{\circ}|=|R^{\circ}|$, $n:=|V(U^{\circ})|=2n_{0}$, $s:=|A^{\circ}|$, and let $\delta,\Delta$ denote the minimum/maximum degrees of $U^{\circ}$; then $\delta=\Theta(s)$ and $\Delta\leq s$.
Mean control in the dense regime.
The remaining quantitative hypothesis is the smallness of the mean-bad probability $p_{\mathrm{mean}}(u)$ from (18). For the dense core from Section 4.7 we have $\delta=\Theta(s)$ and $\Delta\leq s$, hence
$m_{\min}=\alpha+2\lambda\delta=\Theta(\Delta)\qquad\text{and}\qquad\frac{\Delta}{m_{\min}}=\Theta(1).$
Intuitively, the external-drift conditional mean is a smooth (indeed Lipschitz) function of the boundary data, and in the dense regime it concentrates at a constant scale. The next lemma makes this quantitative.
Lemma 4.25 (Mean functional concentration in the dense regime).
Fix $i,t,e=(\ell,r)$ and $d_{0}\in\{\pm\gamma/2\}$, and work in the open-edge regime $\mathsf{Open}_{e,t}$. Define
$M_{e,t}(d_{0}):=\operatorname{E}\!\bigl[R^{\prime}_{e,t}(d_{0})\mid\mathcal{F},f(\ell),f(r)\bigr],\qquad M:=M_{e,t}(d_{0}).$
Let $\mathcal{Y}_{e,t}$ denote the collection of real-valued coordinates among $(\mathcal{F},f(\ell),f(r))$ that remain random after conditioning on $I=i,T_{<t},\mathsf{Open}_{e,t}$; equivalently,
$\mathcal{Y}_{e,t}=\mathcal{F}\cup\{f(u):u\in\{\ell,r\}\ \text{and $f(u)$ is \emph{not} revealed by }T_{<t}\}.$
Recall $m_{\min}:=\alpha+2\lambda\delta$. Under the conditional law given $I=i,T_{<t},\mathsf{Open}_{e,t}$:
  1. 1.
    $M$ is $L_{\mathrm{mean}}$–Lipschitz as a function of the random coordinates $\mathcal{Y}_{e,t}$, with
    $L_{\mathrm{mean}}\leq C_{0}(\lambda+\beta)^{2}\frac{\Delta}{m_{\min}}.$
  2. 2.
    Consequently, for all $u>0$,
    $\Pr\!\left(\left.|M-\operatorname{E}[M]|\geq u/4\right|\,I=i,T_{<t},\ \mathsf{Open}_{e,t}\right)\leq 2\exp\!\left(-\frac{\alpha u^{2}}{32L_{\mathrm{mean}}^{2}}\right).$
    In particular, if $u\geq 4|\operatorname{E}[M]|$ then
    $\Pr\!\left(\left.|M|\geq u/2\right|\,I=i,T_{<t},\ \mathsf{Open}_{e,t}\right)\leq 2\exp\!\left(-\frac{\alpha u^{2}}{32L_{\mathrm{mean}}^{2}}\right).$
Proof.
We keep the notation $\Gamma=N(\ell)\cup N(r)\setminus\{\ell,r\}$. As in Section 4.2,
$R^{\prime}_{e,t}(d_{0})=\sum_{v\in\Gamma}G_{v}\bigl(f(v)\bigr),\qquad\operatorname{Lip}(G_{v})\leq C(\lambda+\beta),$
where each $G_{v}$ is a deterministic one-variable function of $f(v)$ once the discrete data $(i,t,e,d_{0})$ are fixed and we condition on $(f(\ell),f(r))$. In the matching-edge case, $G_{v}$ is in fact independent of $(f(\ell),f(r))$ due to the midpoint integration in Section 4.2. In the open-edge case, $G_{v}$ may depend parametrically on the already revealed endpoint value, but it is independent of every boundary coordinate that remains random under the conditioning (in particular, it does not depend on the unrevealed endpoint variable). Thus, conditioning on $(\mathcal{F},f(\ell),f(r))$, the random variables $\{f(v)\}_{v\in\Gamma}$ are independent and only their one-site laws vary with the boundary. For each $v\in\Gamma$, let $\nu_{v}$ be the one-site conditional law of $f(v)$ given $(\mathcal{F},f(\ell),f(r))$, with potential $h_{v}$; it is $m_{v}$-strongly log-concave for some $m_{v}\geq m_{\min}$.
Step 1 (one-site sensitivity). Fix $v\in\Gamma$ and a boundary coordinate $y$ that belongs to $\mathcal{Y}_{e,t}$ and influences the one-site law $\nu_{v}$; in particular, $y\in\mathrm{bdry}_{v}$ or $y\in\{f(\ell),f(r)\}$ is an unrevealed endpoint variable. Write
$m_{v}(\mathrm{bdry}_{v}):=\operatorname{E}_{\nu_{v}}\bigl[G_{v}(X)\bigr],\qquad X\sim\nu_{v}.$
Differentiating under the integral sign gives
$\frac{\partial}{\partial y}m_{v}(\mathrm{bdry}_{v})=\operatorname{E}_{\nu_{v}}\!\bigl[\partial_{y}G_{v}(X)\bigr]-\operatorname{Cov}_{\nu_{v}}\!\bigl(G_{v}(X),\,\partial_{y}h_{v}(X)\bigr).$
By the discussion above, for every boundary coordinate $y$ that remains random under the conditioning (namely $y\in\mathrm{bdry}_{v}$ or $y$ is an unrevealed endpoint variable), $G_{v}$ has no dependence on $y$. Hence $\partial_{y}G_{v}\equiv 0$ and therefore
$\frac{\partial}{\partial y}m_{v}(\mathrm{bdry}_{v})=-\operatorname{Cov}_{\nu_{v}}\!\bigl(G_{v}(X),\,\partial_{y}h_{v}(X)\bigr).$
Moreover $\operatorname{Lip}(\partial_{y}h_{v})\leq C(\lambda+\beta)$. By the Brascamp–Lieb covariance bound for an $m_{v}$-strongly log-concave measure,
$\left|\frac{\partial}{\partial y}m_{v}(\mathrm{bdry}_{v})\right|\leq\frac{C(\lambda+\beta)^{2}}{m_{v}}\leq\frac{C(\lambda+\beta)^{2}}{m_{\min}}.$
Since $|\mathrm{bdry}_{v}|\leq\Delta$ and there are at most two distinguished boundary coordinates (the endpoints), this implies
$\|\nabla_{\mathrm{bdry}_{v}}m_{v}\|_{2}\leq\sqrt{\Delta}\,\frac{C(\lambda+\beta)^{2}}{m_{\min}},\qquad\text{and for each }y\in\{f(\ell),f(r)\}\cap\mathcal{Y}_{e,t},\ \left|\frac{\partial}{\partial y}m_{v}\right|\leq\frac{C(\lambda+\beta)^{2}}{m_{\min}}.$
Step 2 (global Lipschitzness of $M$). Summing the above over $v\in\Gamma$ and using that the sets $\{\mathrm{bdry}_{v}\}_{v\in\Gamma}$ are disjoint in a $C_{4}$-free graph, we obtain
$\|\nabla_{\mathcal{F}}M\|_{2}\leq\sqrt{\sum_{v\in\Gamma}\|\nabla_{\mathrm{bdry}_{v}}m_{v}\|_{2}^{2}}\leq\sqrt{|\Gamma|\Delta}\,\frac{C(\lambda+\beta)^{2}}{m_{\min}}\leq C(\lambda+\beta)^{2}\frac{\Delta}{m_{\min}}.$
Likewise, if an endpoint coordinate $y\in\{f(\ell),f(r)\}$ is not revealed by $T_{<t}$ (equivalently, $y\in\mathcal{Y}_{e,t}$), then Step 1 gives $|\partial_{y}m_{v}|\leq C(\lambda+\beta)^{2}/m_{\min}$ and hence
$\left|\frac{\partial M}{\partial y}\right|\leq\sum_{v\in\Gamma}\left|\frac{\partial m_{v}}{\partial y}\right|\leq|\Gamma|\frac{C(\lambda+\beta)^{2}}{m_{\min}}\leq C(\lambda+\beta)^{2}\frac{\Delta}{m_{\min}}.$
Since $|\{f(\ell),f(r)\}\cap\mathcal{Y}_{e,t}|\leq 2$, combining this with the bound on $\|\nabla_{\mathcal{F}}M\|_{2}$ yields
$\|\nabla_{\mathcal{Y}_{e,t}}M\|_{2}\leq C_{0}(\lambda+\beta)^{2}\frac{\Delta}{m_{\min}}.$
Thus $M$ is $L_{\mathrm{mean}}$-Lipschitz as a function of $\mathcal{Y}_{e,t}$ with the stated bound, proving (i).
Step 3 (concentration). Conditioning on $I=i,T_{<t},\mathsf{Open}_{e,t}$ fixes some coordinates but preserves $\alpha$-strong log-concavity of the remaining density (the pinning term contributes $\alpha I$ on the free coordinates). Standard Gaussian concentration for $\alpha$-strongly log-concave measures (Bakry–Émery/Herbst) therefore implies that any $L$-Lipschitz function $F$ of the remaining coordinates satisfies
$\Pr\!\left(\left.|F-\operatorname{E}[F]|\geq t\right|I=i,T_{<t},\ \mathsf{Open}_{e,t}\right)\leq 2\exp\!\left(-\frac{\alpha t^{2}}{2L^{2}}\right).$
Applying this with $F=M$ (viewed as a function of $\mathcal{Y}_{e,t}$) and $t=u/4$ gives (ii). The final display follows by the triangle inequality. ∎
Lemma 4.26 (Mean control closes automatically in the dense almost-regular regime).
Consider the parameter choice in Theorem 4.27 (in particular $\alpha=\lambda=1$, $\beta=10$, $B=4\gamma$, $u=\beta\gamma/10$), and assume the base graph $G$ is $C_{4}$-free with degrees satisfying $\delta\leq\deg(v)\leq\Delta$ for all $v$, where $\Delta/\delta=O(1)$ and $\Delta=\Theta(\sqrt{n})$. Then, for all sufficiently large $n$ (equivalently, all sufficiently large $Q=\Theta(\sqrt{n})$), the mean-control quantity from (18) satisfies
$p_{\mathrm{mean}}(u)\leq\frac{1}{400Q^{2}}$
provided the constant $C_{\gamma}$ in the definition of $\gamma$ is chosen large enough.
Proof.
Fix any $i,t,e$ and set
$\mathcal{C}:=\{I=i,\ T_{<t},\ \mathsf{Open}_{e,t}\}.$
By Section 4.7(ii), for each $d_{0}\in\{\pm\gamma/2\}$,
$\Pr\!\left(\left.|M_{e,t}(d_{0})-\operatorname{E}[M_{e,t}(d_{0})]|\geq\frac{u}{4}\right|\mathcal{C}\right)\leq 2\exp\!\left(-\frac{\alpha u^{2}}{32L_{\mathrm{mean}}^{2}}\right).$
In this dense almost-regular core, $\Delta/\delta=O(1)$ and $m_{\min}=\alpha+2\lambda\delta=\Theta(\Delta)$, hence $\Delta/m_{\min}=\Theta(1)$. Therefore Section 4.7(i) gives $L_{\mathrm{mean}}=O(1)$. Thus the right-hand side is $\exp(-\Omega(u^{2}))=\exp(-\Omega(\gamma^{2}))=(100Q^{4})^{-\Omega(C_{\gamma})}$, which is at most $1/(800Q^{2})$ for $C_{\gamma}$ large.
We now show that the centering term $|\operatorname{E}[M_{e,t}(d_{0})]|$ is negligible compared to $u$ in this dense regime, via (26).
Fix $i,t,e$ and $d_{0}\in\{\pm\gamma/2\}$ and condition on a transcript prefix $T_{<t}$ for which $\mathsf{Open}_{e,t}$ holds. Write $S:=\{\ell,r\}\cup\Gamma$ for the one-neighborhood of $e$, and let $\mathcal{F}_{S}$ denote the $\sigma$-field generated by all spins outside $S$. Under $\mu_{i}(\cdot\mid I=i,T_{<t},\mathcal{F}_{S})$ the spins in $S$ have a log-concave density on the box $B,B]^{S}$ with Hamiltonian $H_{S}$ obtained by freezing the exterior configuration. Let $H^{\mathrm{inc}}_{e,t}(x_{\ell},x_{r},x_{\Gamma})$ be the sub-sum of $H_{S}$ consisting of terms from edges incident to $\ell$ or $r$, excluding the designated edge $e$. Define
$V(x_{\ell},x_{r}):=-\log\!\int_{[-B,B]^{\Gamma}}\exp\!\bigl(-H^{\mathrm{inc}}_{e,t}(x_{\ell},x_{r},x_{\Gamma})\bigr)\,dx_{\Gamma}.$
This is the effective (convex) endpoint potential generated by exactly the incident-edge terms that form the drift part $R_{e,t}$ in the one-dimensional decomposition of $W_{e,t}(d)$. (The remaining local terms, i.e., endpoint pinning and the $e$-edge terms, belong to $U_{e,t}$ and $\Psi_{e,i}$ in Section 4.2 and are therefore not part of $R_{e,t}$.) Since we work throughout under $\mathsf{Open}_{e,t}$, at most one endpoint value can already be revealed by the prefix $T_{<t}$. If neither endpoint is revealed then the endpoint marginal is two-dimensional on $B,B]^{2}$; if exactly one endpoint is revealed then the remaining endpoint has a one-dimensional conditional marginal supported on a slice of $B,B]^{2}$. All differentiation and boundary-leakage arguments below are performed after conditioning on any revealed endpoint value, and therefore apply in either the 2D or 1D situation. The only place where the distinction matters is in the definition of the active-strip event $\mathcal{B}_{e,t}$ in (25), where we omit any endpoint that is already fixed by $T_{<t}$.
A direct differentiation under the integral (Leibniz rule) identifies the conditional-mean drift as the derivative of an external effective endpoint potential in the same one-dimensional direction that defines the difference coordinate $d$ in Section 4.2. Under $\mathsf{Open}_{e,t}$ there are three cases, and in each case we parametrize the endpoints as a path $d\mapsto(x_{\ell}(d),x_{r}(d))$ so that $R_{e,t}(d)$ is obtained by evaluating the incident edge terms at $(x_{\ell}(d),x_{r}(d))$:
  • If neither endpoint is revealed in $T_{<t}$, we work at fixed midpoint $z:=\tfrac{1}{2}(f(\ell)+f(r))$ and set $x_{\ell}(d)=z+d/2$ and $x_{r}(d)=z-d/2$.
  • If $f(\ell)$ is revealed (but $f(r)$ is not), we set $a:=f(\ell)$ and vary only the right endpoint: $x_{\ell}(d)=a$ and $x_{r}(d)=a-d$.
  • If $f(r)$ is revealed (but $f(\ell)$ is not), we set $b:=f(r)$ and vary only the left endpoint: $x_{\ell}(d)=b+d$ and $x_{r}(d)=b$.
In all three cases, differentiating under the integral and applying the chain rule yields, for every $d_{0}\in\{\pm\gamma/2\}$,
$M_{e,t}(d_{0})\;=\;\left.\frac{d}{dd}\,V\!\bigl(x_{\ell}(d),x_{r}(d)\bigr)\right|_{d=d_{0}}.$ (23)
Equivalently,
[…]
In particular, the evaluation point depends on $d_{0}$ (and on which endpoint, if any, is fixed), so (23) does not imply $M_{e,t}(+\gamma/2)=M_{e,t}(-\gamma/2)$.
To bound the mean-shift term $|\operatorname{E}[M_{e,t}(d_{0})]|$, we isolate the “box effect” of the hard truncation. In the untruncated model (no indicator on [B,B][-B,B]) the same differentiation-under-the-integral calculation that gives (23) yields a perfectly centered score identity (an integration-by-parts/score-centering argument; see Section 4.1), and in particular the corresponding mean term vanishes. Therefore $\operatorname{E}[M_{e,t}(d_{0})]$ is entirely due to boundary contributions coming from the hard box.
Let
$\eta_{\mathrm{strip}}:=C_{\mathrm{bdry}}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}\cdot\exp\!\left(-c_{\mathrm{bdry}}\cdot\delta\cdot\frac{(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2}\right),$ (24)
where $c_{\mathrm{bdry}},C_{\mathrm{bdry}}>0$ are the absolute constants from Section 4.1. For $d_{0}\in\{\pm\gamma/2\}$, varying the difference coordinate $d$ moves the endpoint pair $(x_{\ell}(d),x_{r}(d))$ along the corresponding path discussed above (a diagonal line in the 2D case, and a horizontal/vertical line in the 1D cases) within the box $B,B]^{2}$. When differentiating the truncated integral defining $V(x_{\ell}(d),x_{r}(d))$ along this path, boundary contributions can arise only when the path hits a face of the box. Since $R^{\prime}_{e,t}(d)$ is the derivative with respect to $d$ in the positive direction, the only face that can contribute with the adverse sign for the relevant one-sided mean control is the face approached when $d$ increases. In all cases of $\mathsf{Open}_{e,t}$ this active face is always one of the two oriented faces $x_{\ell}=+B$ (whenever $x_{\ell}$ increases with $d$) and $x_{r}=-B$ (whenever $x_{r}$ decreases with $d$), independently of the sign of $d_{0}$.
Because a transcript prefix may already reveal one endpoint value, we define the corresponding active-strip event by omitting any endpoint that is fixed by $T_{<t}$:
[…]
On $\mathcal{B}_{e,t}^{c}$, the probe segment stays a distance $\geq\gamma/2$ away from the active boundary faces, so the boundary terms vanish and the remaining interior term cancels by the (untruncated) score centering (cf. Section 4.1). Thus, using the uniform bound $|M_{e,t}(d_{0})|\leq\mathrm{poly}(\Delta,\lambda,\beta,B,\gamma)$ (a sum of $O(\Delta)$ one-edge forces, each of size $O((\lambda+\beta)(B+\gamma))$), we obtain
$|\operatorname{E}[M_{e,t}(d_{0})]|\ \leq\ \mathrm{poly}(\Delta,\lambda,\beta,B,\gamma)\cdot\Pr[\mathcal{B}_{e,t}].$
Finally, Section 4.1 applies uniformly under any boundary conditioning, so whichever endpoint remains random under $T_{<t}$ satisfies the corresponding strip bound. From the definition (25) of $\mathcal{B}_{e,t}$ we therefore have, uniformly over all transcript prefixes for which $\mathsf{Open}_{e,t}$ holds,
$\Pr[\mathcal{B}_{e,t}]\leq 2\eta_{\mathrm{strip}}.$
Combining gives
$|\operatorname{E}[M_{e,t}(d_{0})]|\;\leq\;\mathrm{poly}(\Delta,\lambda,\beta,B,\gamma)\cdot\eta_{\mathrm{strip}},$ (26)
as claimed.
In the present dense regime, Section 4.1 gives $\eta_{\mathrm{strip}}\leq\exp(-\Omega(\delta\gamma^{2}))$ with $\delta=\Theta(\Delta)=\Theta(\sqrt{n})$, so the right-hand side of (26) is superpolynomially small in $Q$. In particular, for all large $n$ it is at most $u/4$.
Combining this with the deviation bound above yields
$\Pr\!\left(\left.|M_{e,t}(d_{0})|\geq\frac{u}{2}\right|I=i,T_{<t},\ \mathsf{Open}_{e,t}\right)$ $\leq\Pr\!\left(\left.|M_{e,t}(d_{0})-\operatorname{E}[M_{e,t}(d_{0})]|\geq\frac{u}{4}\right|I=i,T_{<t},\ \mathsf{Open}_{e,t}\right)$
$\leq\frac{1}{800Q^{2}}.$
Taking a union bound over $d_{0}\in\{\pm\gamma/2\}$ and then the supremum over $i,t,e$ (as well as transcript prefixes with $\mathsf{Open}_{e,t}$) gives $p_{\mathrm{mean}}(u)\leq 1/(400Q^{2})$. ∎
Theorem 4.27 (Deterministic adaptive one-sided lower bound from the truncated Gibbs model).
There exist absolute constants $c,\varepsilon>0$ such that for infinitely many $n$ there is a bipartite DAG $G$ on $n$ vertices with the following property. Every deterministic adaptive one-sided $\varepsilon$-tester for monotonicity on $G$ requires at least $c\sqrt{n}$ queries.
Proof.
Apply Section 4.7 with the dense choice $P=\lfloor c_{P}N\rfloor$ for a sufficiently small absolute constant $c_{P}\in(0,1/10)$. Let $G:=G^{\circ}$ be the resulting bipartite DAG together with the family of edge-disjoint positive matchings $\{M_{i}\}_{i\in[s]}$ (indexed by $A^{\circ}$ in Section 4.7). Write $n_{0}:=|L^{\circ}|=|R^{\circ}|$ and $n:=|V(G)|=2n_{0}$. In particular,
$s=\Theta(\sqrt{n_{0}})=\Theta(\sqrt{n}),\qquad|M_{i}|\geq\varepsilon_{0}n_{0}\ \ (\forall i\in[s]),\qquad\text{and}\qquad c_{\delta}s\leq\deg(v)\leq s\ \ (\forall v),$
so $\delta:=\min_{v}\deg(v)=\Theta(\sqrt{n_{0}})=\Theta(\sqrt{n})$ and $\Delta:=\max_{v}\deg(v)=\Theta(\sqrt{n})$ with $\Delta/\delta=O(1)$.
Set
$Q:=\left\lfloor c_{Q}\sqrt{n_{0}}\right\rfloor$
for a sufficiently small absolute constant $c_{Q}>0$ (chosen at the end), and instantiate the truncated Gibbs family $\{\mu_{i}\}_{i\in[s]}$ on $G$ with parameters
$\alpha=1,\qquad\lambda=1,\qquad\beta=10,\qquad\gamma:=\sqrt{C_{\gamma}\log(100Q^{4})},\qquad B:=4\gamma,\qquad u:=\frac{\beta\gamma}{10},$
where $C_{\gamma}$ is a sufficiently large absolute constant.
We verify the conditions of Theorem 4.22.
(i) Support and separation. The support condition (15) holds since $B=4\gamma$. For the separation condition (16), the right-hand side equals
$\alpha\!\left(B+\frac{\gamma}{2}\right)+\lambda\gamma+u\;=\;\left(4+\frac{1}{2}+1+\frac{\beta}{10}\right)\gamma\;=\;6.5\,\gamma,$
while the left-hand side is $\beta\gamma=10\gamma$, so (16) holds.
(ii) Tail probabilities $p_{\mathrm{match}}$ and $p_{\mathrm{spur}}$. With $\alpha=\lambda=1$ we have $m_{\mathrm{edge}}=\alpha/2+2\lambda=5/2$, so
$\exp\!\left(-\frac{m_{\mathrm{edge}}\gamma^{2}}{8}\right)=\exp\!\left(-\frac{5}{16}\gamma^{2}\right)=(100Q^{4})^{-5C_{\gamma}/16}.$
Choosing $C_{\gamma}$ large makes this at most $\min\{10^{-4},\,1/(400Q^{2})\}$.
For the drift term, note that $m_{\min}=\alpha+2\lambda\delta=\Theta(\sqrt{n})$ and $\Delta=\Theta(\sqrt{n})$, hence $m_{\min}/\Delta=\Theta(1)$. With $u=\beta\gamma/10$ the drift tail (12) becomes $p_{\mathrm{drift}}(u/2)\leq 2\exp(-\Omega(\gamma^{2}))$, so for $C_{\gamma}$ large we have $p_{\mathrm{drift}}(u/2)\leq 1/(400Q^{2})$. Finally, Section 4.7 supplies the required mean-control bound $p_{\mathrm{mean}}(u)\leq 1/(400Q^{2})$ for this instantiation. Combining these with (20) gives $p_{\mathrm{match}}\leq 10^{-4}$ and $p_{\mathrm{spur}}\leq 1/(100Q^{2})$.
(iii) Boundary-margin failures. Section 4.3 bounds the boundary failure probability by
$\eta_{0}\;\leq\;Q\cdot C_{\mathrm{bdry}}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}\cdot\exp\!\left(-c_{\mathrm{bdry}}\cdot\frac{\delta(\beta-\lambda)^{2}}{\lambda+\beta}\cdot\gamma^{2}\right).$
Here $\delta=\Theta(\sqrt{n})$ and $\gamma^{2}=\Theta(\log Q)$. Moreover, under our instantiation $B=4\gamma$ and $\lambda,\beta$ are fixed absolute constants, so the prefactor $C_{\mathrm{bdry}}\cdot\frac{\lambda+\beta}{\beta-\lambda}\cdot\frac{B+\gamma}{\gamma}$ is an absolute constant. Thus the exponent dominates, so the exponent is $\Theta(\sqrt{n}\log Q)$ and in particular $\eta_{0}\leq 0.01$ for all large $n$.
(iv) Leakage. Since $m_{\min}=\Theta(\sqrt{n})$ and $B+\gamma=\Theta(\gamma)$, Section 4.4 gives
$\kappa_{Q}\;\leq\;\frac{C_{b}}{\alpha}\cdot\frac{\beta^{2}(B+\gamma)^{2}}{m_{\min}}\;=\;O\!\left(\frac{\gamma^{2}}{\sqrt{n}}\right)\;=\;O\!\left(\frac{\log Q}{\sqrt{n}}\right).$
Thus $Q\kappa_{Q}=O(c_{Q}\log Q)$.
(v) Apply the fixed-budget lower bound. Theorem 4.22 states that any deterministic adaptive one-sided tester with query budget $q\leq Q$ must satisfy
$q\,\kappa_{Q}\ \geq\ (0.65-\eta_{0})\log s-1.$
Since $\eta_{0}\leq 0.01$ and $s=\Theta(\sqrt{n})$ we have $\log s=\Theta(\log Q)$. On the other hand, $Q\kappa_{Q}=O(c_{Q}\log Q)$. Choosing the absolute constant $c_{Q}>0$ small enough makes $Q\kappa_{Q}<(0.65-\eta_{0})\log s-1$ for all large $n$, so no adaptive one-sided tester using at most $Q$ queries can succeed. This yields the stated $\Omega(\sqrt{n})$ lower bound. ∎
Theorem 4.28 (Randomized adaptive one-sided lower bound).
There exist absolute constants $c,\varepsilon>0$ such that for infinitely many $n$ there is a bipartite DAG $G$ on $n$ vertices with the following property. Every randomized adaptive one-sided $\varepsilon$-tester for monotonicity on $G$ requires at least $c\sqrt{n}$ queries.
Proof.
Apply Yao’s minimax principle to the instance distribution induced by sampling $I\sim{\rm Unif}[s]$ and then $f\sim\mu_{I}$. Theorem 4.27 gives the corresponding lower bound for deterministic adaptive one-sided testers under this distribution, hence it also bounds randomized testers. ∎
5 Bounds in terms of $m$ and $\ell$
This section collects bounds parameterized by the edge counts of the transitive reduction and closure. The first two subsections give non-adaptive one-sided testers with self-contained proofs. The final subsection gives a non-adaptive two-sided lower bound in terms of $m$.
Model.
Since the graph is given explicitly and we count only queries to $f$ (see Section 2), we may precompute $\mathrm{TC}(G)$ or perform reachability tests and sample uniformly from $E(\mathrm{TC}(G))$ without affecting query complexity. In particular, the sampling from $E(\mathrm{TC}(G))$ in Algorithm 1 and the reachability checks in Algorithm 2 are treated as free.
5.1 A $O(\sqrt{m\ell}/(\varepsilon n))$-query tester via the transitive reduction
In this subsection, we give a simple one-sided error tester whose query complexity depends on the number of edges in the transitive closure and the number of edges in the transitive reduction.
Throughout this subsection we assume that the input DAG is already transitively reduced and write $m:=|E|$ (see Section 2).
Theorem 5.1.
There exists a one-sided error monotonicity tester with query complexity
$O\!\left(\frac{\sqrt{m\,\ell}}{\varepsilon n}\right),$
where $\ell=|E(\mathrm{TC}(G))|$.
[…]
2   Set $\ell\leftarrow|E(\mathrm{TC}(G))|$;
3   Set $q\leftarrow\Theta\!\left(\frac{\sqrt{m\,\ell}}{\varepsilon n}\right)$;
4 for $i\leftarrow 1$ to $q$ do
5      Sample an edge $(u,v)\in E$ uniformly at random;
6      Query $f(u)$ and $f(v)$;
7    if $f(u)>f(v)$ then reject;
8
9 for $i\leftarrow 1$ to $q$ do
10      Sample a pair $(u,v)\in E(\mathrm{TC}(G))$ uniformly at random;
11      Query $f(u)$ and $f(v)$;
12    if $f(u)>f(v)$ then reject;
13
14  Accept;
15
Algorithm 1 Monotonicity Tester via Transitive Reduction
Proof of Theorem 5.1.
The tester is one-sided error because it rejects only after finding a violating pair.
Fix a function $f:V\to\mathbb{R}$ that is $\varepsilon$-far from monotonicity. By the standard matching characterization of the distance to monotonicity (via Kőnig’s theorem), there exists a matching $M$ of violating pairs in $\mathrm{TC}(G)$ with
$|M|\geq\varepsilon n.$
Let
$F_{f}:=\{(x,y)\in E:f(x)>f(y)\}$
be the set of violating edges in $G$.
We use the following dichotomy, governed by a threshold parameter $\tau>0$ to be chosen later.
Case 1: $|F_{f}|\geq\tau$.
In the first loop of Algorithm 1, each sampled edge of $E$ is violating with probability at least $\tau/m$. Thus, by choosing the hidden constant in $q=\Theta(m/\tau)$ large enough, the rejection probability in the first loop is at least
$1-\left(1-\frac{\tau}{m}\right)^{q}\geq\frac{2}{3}.$
Case 2: $|F_{f}|<\tau$.
For each violating pair $(u,v)\in M$, choose an arbitrary directed path $P_{uv}$ from $u$ to $v$ in $G$. Along $P_{uv}$, since $f(u)>f(v)$, there must exist at least one edge $(x,y)\in E$ with $f(x)>f(y)$. Let $\psi(u,v)$ be the first such violating edge on the path $P_{uv}$, and partition
$M=\biguplus_{e\in F_{f}}M_{e},\qquad M_{e}:=\{(u,v)\in M:\psi(u,v)=e\}.$
For each $e\in F_{f}$, let $A_{e}$ (resp., $B_{e}$) be the set of left (resp., right) endpoints of pairs in $M_{e}$. Since $M$ is a matching, the families $\{A_{e}\}_{e}$ and $\{B_{e}\}_{e}$ are pairwise disjoint, and $|A_{e}|=|B_{e}|=|M_{e}|$.
Fix $e=(x,y)\in F_{f}$. Let $t_{e}$ be a median of the multiset $\{f(a):(a,b)\in M_{e}\}$ and define
$A_{e}^{+}\ :=\ \{a\in A_{e}:\ f(a)\geq t_{e}\},\qquad B_{e}^{-}\ :=\ \{b\in B_{e}:\ f(b)<t_{e}\}.$
Then $|A_{e}^{+}|,|B_{e}^{-}|\geq|M_{e}|/2$ (again because each $(a,b)\in M_{e}$ has $f(a)>f(b)$). For any $a\in A_{e}^{+}$ and $b\in B_{e}^{-}$ we claim that $(a,b)\in E(\mathrm{TC}(G))$ and it is a violating pair. Indeed, by construction of the partition, $a$ reaches $x$ in $G$ along $P_{a,\cdot}$, and $y$ reaches $b$ in $G$ along $P_{\cdot,b}$. Concatenating these paths with the edge $(x,y)$ yields a directed path from $a$ to $b$ in $G$, hence $(a,b)\in E(\mathrm{TC}(G))$, and $f(a)>f(b)$. Therefore, for each $e$ there are at least $|A_{e}^{+}|\cdot|B_{e}^{-}|\geq|M_{e}|^{2}/4$ violating pairs in $\mathrm{TC}(G)$ witnessed through $e$. Summing over $e$ and applying Cauchy–Schwarz gives that the total number of violating pairs in $\mathrm{TC}(G)$ is at least
$\frac{1}{4}\sum_{e\in F_{f}}|M_{e}|^{2}\ \geq\ \frac{1}{4}\cdot\frac{\left(\sum_{e\in F_{f}}|M_{e}|\right)^{2}}{|F_{f}|}\ =\ \frac{|M|^{2}}{4|F_{f}|}\ >\ \frac{\varepsilon^{2}n^{2}}{4\tau}.$
Hence, in the second loop of Algorithm 1, a uniformly sampled edge of $E(\mathrm{TC}(G))$ is violating with probability at least $\varepsilon^{2}n^{2}/(4\tau\ell)$. By choosing the hidden constant in $q=\Theta(\tau\ell/(\varepsilon^{2}n^{2}))$ large enough, the rejection probability in the second loop is at least
$1-\left(1-\frac{\varepsilon^{2}n^{2}}{4\tau\ell}\right)^{q}\geq\frac{2}{3}.$
It remains to balance the two cases. Setting
$\tau:=\varepsilon n\sqrt{\frac{m}{\ell}}$
makes $m/\tau$ and $\tau\ell/(\varepsilon^{2}n^{2})$ equal up to constants, and thus Algorithm 1 has query complexity
$q=\Theta\!\left(\frac{\sqrt{m\,\ell}}{\varepsilon n}\right).$
This completes the proof. ∎
5.2 A simple $O(m^{1/3}/\varepsilon^{2/3})$-query tester
In this subsection, we describe a simple one-sided error tester whose query complexity depends only on the number of edges in the input DAG and $\varepsilon$. Throughout this subsection, we assume that the input DAG $G=(V,E)$ is transitively reduced and write $m:=|E|$ (see Section 2).
[…]
2 $q\leftarrow\Theta\!\left(m^{1/3}/\varepsilon^{2/3}\right)$;
3   Sample $q$ edges $(u,v)\in E$ uniformly at random;
4   Query $f(u),f(v)$ for each sampled edge; if any sampled edge is violating, reject;
5   Sample two multisets $Q_{L},Q_{R}$ of $q$ vertices uniformly at random with replacement;
6   Query $f(v)$ for each $v\in Q_{L}\cup Q_{R}$ (reusing answers for repeated vertices);
7 if there exist $x\in Q_{L}$ and $y\in Q_{R}$ such that $(x,y)\in E(\mathrm{TC}(G))$ and $f(x)>f(y)$ then reject;
8   Accept;
9
Algorithm 2 Two-stage monotonicity tester
Theorem 5.2.
Algorithm 2 is a one-sided error monotonicity tester with query complexity
$O\!\left(m^{1/3}/\varepsilon^{2/3}\right).$
Proof.
It is immediate that Algorithm 2 has one-sided error: if $f$ is monotone, then there are no violating pairs in $\mathrm{TC}(G)$, and in particular there are no violating edges in $E$, so the algorithm always accepts.
Fix a function $f:V\to\mathbb{R}$ that is $\varepsilon$-far from monotonicity. By the standard matching characterization of the distance to monotonicity (via Kőnig’s theorem), there exists a matching $M$ of violating pairs in $\mathrm{TC}(G)$ such that
$|M|\geq\varepsilon n.$ (27)
Let
$F_{f}:=\{(x,y)\in E:f(x)>f(y)\}$
be the set of violating edges in $G$.
We will use a dichotomy based on the size of $F_{f}$. Set the threshold
$\tau:=(\varepsilon m)^{2/3}.$
We show that with $q=\Theta(m^{1/3}/\varepsilon^{2/3})$, the algorithm rejects with probability at least $2/3$ in either case.
Case 1: $|F_{f}|\geq\tau$.
In the first stage of Algorithm 2, each sampled edge is violating with probability at least $|F_{f}|/m\geq\tau/m$. Hence,
$\Pr[\text{fail to reject in stage 1}]\leq\left(1-\frac{\tau}{m}\right)^{q}\leq\exp\!\left(-q\frac{\tau}{m}\right).$
Choosing the hidden constant in $q=\Theta(m/\tau)=\Theta(m^{1/3}/\varepsilon^{2/3})$ sufficiently large, the right-hand side becomes at most $1/3$, so stage 1 rejects with probability at least $2/3$.
Case 2: $|F_{f}|<\tau$.
We analyze the second stage using the standard “bottleneck partition” and the Poisson approximation argument.
For each violating pair $(u,v)\in M$, fix an arbitrary directed path $P_{uv}$ from $u$ to $v$ in $G$. Since $f(u)>f(v)$, along $P_{uv}$ there must exist at least one edge $(x,y)\in E$ with $f(x)>f(y)$; define $\psi(u,v)$ to be the first such edge on $P_{uv}$. This yields a partition
$M=\biguplus_{e\in F_{f}}M_{e},\qquad M_{e}:=\{(u,v)\in M:\ \psi(u,v)=e\}.$
For each $e\in F_{f}$, let $A_{e}$ be the set of left endpoints of pairs in $M_{e}$ and let $B_{e}$ be the set of right endpoints. Since $M$ is a matching, the families $\{A_{e}\}_{e\in F_{f}}$ and $\{B_{e}\}_{e\in F_{f}}$ are pairwise disjoint, and $|A_{e}|=|B_{e}|=|M_{e}|$.
Fix $e=(x,y)\in F_{f}$. Let $t_{e}$ be a median of the multiset $\{f(a):(a,b)\in M_{e}\}$ and define
$A_{e}^{+}\ :=\ \{a\in A_{e}:\ f(a)\geq t_{e}\},\qquad B_{e}^{-}\ :=\ \{b\in B_{e}:\ f(b)<t_{e}\}.$
Then $|A_{e}^{+}|,|B_{e}^{-}|\geq|M_{e}|/2$ (since each $(a,b)\in M_{e}$ has $f(a)>f(b)$). For any $a\in A_{e}^{+}$ and $b\in B_{e}^{-}$, we have $a\leadsto x$ and $y\leadsto b$ in $G$ by construction of the “first violating edge” assignment, hence $a\leadsto b$ in $G$ via the path $a\leadsto x\to y\leadsto b$. Therefore $(a,b)\in E(\mathrm{TC}(G))$ and $(a,b)$ is a violating pair. Consequently, if the sampled multiset $Q_{L}$ contains at least one vertex in $A_{e}^{+}$ and $Q_{R}$ contains at least one vertex in $B_{e}^{-}$ for some $e\in F_{f}$, then stage 2 rejects.
Let $X_{v}^{L}$ (resp., $X_{v}^{R}$) be the number of times $v\in V$ is sampled when constructing $Q_{L}$ (resp., $Q_{R}$). We bound the failure event
$\mathcal{E}:=\bigl[\ \forall e\in F_{f},\ \sum_{v\in A_{e}^{+}}X_{v}^{L}=0\ \ \vee\ \ \sum_{v\in B_{e}^{-}}X_{v}^{R}=0\ \bigr].$
Introduce independent Poisson variables $Y_{v}^{L},Y_{v}^{R}\sim\mathrm{Pois}(\lambda)$ with mean $\lambda:=q/n$. By the Poisson approximation theorem [4],
$\Pr[\mathcal{E}]\leq 2\Pr\Bigl[\ \forall e\in F_{f},\ \sum_{v\in A_{e}^{+}}Y_{v}^{L}=0\ \ \vee\ \ \sum_{v\in B_{e}^{-}}Y_{v}^{R}=0\ \Bigr].$ (28)
Since the sets $\{A_{e}^{+}\}_{e}$ are disjoint and the sets $\{B_{e}^{-}\}_{e}$ are disjoint, and the left/right samples are independent, the events inside the probability in (28) are independent, and thus
$\Pr\Bigl[\ \forall e\in F_{f},\ \sum_{v\in A_{e}^{+}}Y_{v}^{L}=0\ \ \vee\ \ \sum_{v\in B_{e}^{-}}Y_{v}^{R}=0\ \Bigr]=\prod_{e\in F_{f}}\left(1-\bigl(1-e^{-\lambda|A_{e}^{+}|}\bigr)\bigl(1-e^{-\lambda|B_{e}^{-}|}\bigr)\right).$
Using $|A_{e}^{+}|,|B_{e}^{-}|\geq|M_{e}|/2$, we can lower bound the success probability for each $e$ by $(1-e^{-\lambda|M_{e}|/2})^{2}$, which only changes constants in the analysis below.
If there exists $e$ with $\lambda|M_{e}|\geq 6$, then $\left(1-e^{-\lambda|M_{e}|/2}\right)^{2}\geq(1-e^{-3})^{2}$ and hence the rejection probability is at least $1-2(1-(1-e^{-3})^{2})>2/3$.
Otherwise, $\lambda|M_{e}|\leq 6$ for all $e\in F_{f}$. For $x\in[0,3]$ we have $1-e^{-x}\geq x/4$, so
$1-\left(1-e^{-\lambda|M_{e}|/2}\right)^{2}\leq 1-\frac{\lambda^{2}|M_{e}|^{2}}{64}.$
Therefore,
$\Pr[\mathcal{E}]\leq 2\exp\!\left(-\frac{\lambda^{2}}{64}\sum_{e\in F_{f}}|M_{e}|^{2}\right).$
By Cauchy–Schwarz and (27),
$\sum_{e\in F_{f}}|M_{e}|^{2}\ \geq\ \frac{\left(\sum_{e\in F_{f}}|M_{e}|\right)^{2}}{|F_{f}|}\ =\ \frac{|M|^{2}}{|F_{f}|}\ \geq\ \frac{\varepsilon^{2}n^{2}}{|F_{f}|}\ >\ \frac{\varepsilon^{2}n^{2}}{\tau},$
where we used the assumption $|F_{f}|<\tau$ in the last inequality. Plugging this in and using $\lambda=q/n$, we obtain
$\Pr[\mathcal{E}]\leq 2\exp\!\left(-\frac{\varepsilon^{2}q^{2}}{64\tau}\right).$
Finally, with $\tau=(\varepsilon m)^{2/3}$ we have $\sqrt{\tau}/\varepsilon=m^{1/3}/\varepsilon^{2/3}$. Thus, by setting $q=\Theta(\sqrt{\tau}/\varepsilon)=\Theta(m^{1/3}/\varepsilon^{2/3})$ with a sufficiently large hidden constant, the above probability becomes at most $1/3$, and stage 2 rejects with probability at least $2/3$.
In both cases, Algorithm 2 rejects $\varepsilon$-far functions with probability at least $2/3$, and it uses $O(q)=O(m^{1/3}/\varepsilon^{2/3})$ queries. ∎
Acknowledgements
We thank Nathan Harms, Jane Lange, Mikhail Makarov, Cameron Seth, and Yubo Zhang for helpful discussions. Y.Y. is supported by JSPS KAKENHI Grant Number 22H05001 and 25K24465.